### v2

# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

Abstract.
The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps). Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice. In this work, we conduct a user-centered analysis of 676,066 reviews from 173 Gen-AI apps on the Google Play Store. We introduce a four-phase methodology, SARA (Selection, Acquisition, Refinement, and Analysis), that enables the systematic extraction of user insights using prompt-based LLM techniques. First, we demonstrate the reliability of LLMs in topic extraction, achieving 91% accuracy through five-shot prompting and non-informative review filtering. Then, we apply this method to the informative reviews, identify the top 10 user-discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities. Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps. Based on our findings and observations, we present actionable implications for developers and researchers.
Generative AI, Large Language Models, User Review Analysis, Mobile Apps
ccs: Software and its engineering Software maintenance tools
1. Introduction
Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod and Agal, 2023). A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days. This rapid adoption signalled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces. In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci, 2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat. Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel et al., 2024; Zhang et al., 2024). Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (Gozalo-Brizuela and Garrido-Merchán, 2023).
Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi et al., 2023). LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue. Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau et al., 2025). When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen and Schlote, 2024). A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs). APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen et al., 2025b).
While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges. Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen et al., 2025a). These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao et al., 2025). Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern. Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar et al., 2024). As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
However, despite growing interest in integrating LLMs into mobile apps, little is known about the specific concerns and experiences of end-users. Existing research primarily focused on the challenges faced by developers. For example, Nahar et al. (Nahar et al., 2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness. While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience. Chen et al. (Chen et al., 2025a) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development. This study offers insight into implementation bottlenecks but also neglects user-facing concerns. Similarly, Shao et al. (Shao et al., 2025) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues. However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice. Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a; Shata and Hartley, 2025; Kim et al., 2025a; Yuen and Schlote, 2024), creative image generation (Tang et al., 2024; Haase et al., 2023; Oppenlaender et al., 2023), or mental health (Jin et al., 2025; Heinz et al., 2025). Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas. To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains. We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration. This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
Figure 1. Examples of reviews of Gen-AI apps
To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps. For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration. Building on this foundation, we introduce a four-phase methodology, Selection, Acquisition, Refinement, and Analysis (SARA). Our proposed methodology enables the large-scale collection and processing of user reviews from Gen-AI apps, supporting effective analysis using LLMs. First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition). We then apply a data cleaning pipeline to enhance the quality of input data for LLM-based analysis (i.e., Refinement). Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis). Using this methodology, we conduct an empirical study on 676,066 reviews from 173 Gen-AI apps to address the following research questions (RQs):
RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
{adjustwidth}
0.8cm In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews. Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy. The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM pipeline for downstream analysis.
RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
{adjustwidth}
0.8cm To answer this RQ, we apply the LLM-based topic labelling pipeline from RQ1 to informative reviews across 173 Gen-AI apps. We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers. Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset. Notably, some users form emotional bonds with conversational agents and value personalized interactions. These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
RQ3: What temporal trends emerge in user feedback on Gen-AI topics? {adjustwidth}0.8cm In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures. We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation. Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time. These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
Our study provides valuable insights for app developers and researchers by uncovering how users perceive and evaluate Gen-AI features in mobile apps. Specifically, our contributions are as follows:
  • We propose a four-phase methodology, SARA, for analyzing user reviews of Gen-AI apps. Applying this methodology, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labelled examples and filtering out non-informative reviews significantly improves accuracy.
  • We provide actionable feedback for developers to maintain competitiveness, for policymakers to explore the ethical concerns surrounding content policy and censorship, and for user requirement engineers to better understand user needs.
  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labelled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
We organize the remainder of this paper as follows. Section 2 illustrates existing work on analyzing user reviews. Section 3 presents our methodology. Sections 46 describe the approach and the results obtained in our RQs. Section 7 synthesizes our key findings and implications of our work. Section 8 discusses potential threats to the validity of our findings. Finally, Section 9 concludes the paper and discusses implications for future research.
2. Related Work
In our literature review, we examine three key areas: 1) the user perspectives of Gen-AI apps, 2) user review analysis techniques, and 3) studies on prompt design and evaluation.
2.1. User Perspective of Gen-AI Apps
The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a). Shata and Hartley (Shata and Hartley, 2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption. Kim et al. (Kim et al., 2025a) compared student and faculty perceptions on the use of Gen-AI in learning. Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn. Both faculty and students believed that Gen-AI will have more negative than positive effects on learning. Yuen and Schlote (Yuen and Schlote, 2024) surveyed adult learners to explore their views on the integration of AI in language learning apps. Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
In the creative image domain, Tang et al. (Tang et al., 2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools. They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs. In support of that, Haase et al. (Haase et al., 2023) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation. Oppenlaender et al. (Oppenlaender et al., 2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey. Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (Sharma and Patel, 2024; Madhuri et al., 2025; S et al., 2025; Kheterpal and Gill, 2024). Our focus is on Gen-AI-driven treatment through conversational agents. Jin et al. (Jin et al., 2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention. Heinz et al. (Heinz et al., 2025) studied the efficacy of a mental health AI chatbot. They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
Unlike prior work that relied on expert synthesis or self-reported perceptions, our study analyzes large-scale, real-world reviews from users actively engaging with Gen-AI apps. We offer a complementary perspective on actual usage and feedback. We also take a broad perspective by examining Gen-AI apps across all domains, without prioritizing any specific application area.
2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews and 2) topic extraction and assignment to user reviews.
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modelling. Ghosh et al. (Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
Topic extraction and assignment to user reviews. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modelling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi et al., 2021a) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. Assi et al. (Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combines Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, aiming to uncover ethically relevant user feedback that traditional keyword-based methods often overlook. Their approach achieved an F1-score of 81%.
Instead of extracting and assigning topics to user reviews, prior studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos Santos et al., 2023) extracted topics from user reviews without performing the assignment task. Their study examined the impact of app updates on accessibility by utilizing GPT-4 to identify key issues, barriers, improvements, update consequences, and user sentiments. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis. Ren et al. (Ren et al., 2024) used LDA for topic extraction and pretrained LLMs for requirement extraction. Rezaei Nasab et al. (Rezaei Nasab et al., 2025) conducted a study of AI-based mobile app reviews to identify fairness-related concerns. They constructed a labelled dataset of fairness vs. non-fairness reviews, developed machine learning and LLM-based classifiers to detect fairness mentions, applied K-Means clustering, and performed manual coding to identify six types of fairness concerns.
Our work builds on this line of research by utilizing LLMs to filter out non-informative reviews and identify topics in user reviews of Gen-AI apps. Based on the analysis of user reviews, we identify the challenges and potential opportunities to adopt Gen-AI functionalities in mobile apps.
2.3. Prompt Design and Evaluation
Prompt engineering is the process of designing and refining prompts to effectively guide LLMs in performing specific tasks (Koyuncu, 2025). Prior work discusses prompt design in three directions: 1) the process of prompt design, 2) the definition of the required tasks, and 3) identifying the number of included examples.
The process of prompt design. Marvin et al. (Marvin et al., 2024) emphasized the importance of prompt refinement and outlined several prompt engineering techniques, including dynamic prompting, which refers to iteratively adjusting the prompt based on the generated output. He et al. (He et al., 2024) investigated the impact of prompt format across various tasks and found that prompt structure significantly affects performance, with different formats performing best in different task types. In our work, we adopt a prompt format inspired by prior work (Assi et al., 2025a) and refine it to suit our study’s goals, employing dynamic prompting to enhance accuracy and relevance further.
The definition of the required tasks. To ensure accurate, ethical, and reliable AI outputs, effective prompt engineering requires understanding common pitfalls, such as ambiguity, bias, and lack of context (Giray, 2023). Prompts can be explicit, implicit, or creative, varying in the amount of guidance they offer to the AI. Explicit prompts provide clear instructions, while implicit and creative prompts allow more flexibility and imagination but may produce less predictable results (Data Science Horizons, 2023). Ho et al. (Ho et al., 2024) adopted a prompt classification framework consisting of the three types. The authors analyzed how different prompts affect AI-generated responses. Explicit prompts generate specific and fact based answers, implicit prompts produce broader and more flexible responses, while creative prompts encourage innovation and originality. In our work, we adopt an explicit prompting strategy to improve structure and reproducibility.
Identifying the number of included examples. Deshmukh et al. (Deshmukh et al., 2025) conducted a literature review of prompt engineering and discussed 16 LLM prompt patterns to enhance output accuracy. Their analysis highlighted the importance of clarity, specificity, defined output formats, and contextual examples to improve model accuracy. However, it did not address techniques for optimizing example selection (i.e., determining the optimal number of examples to include in the prompt). Complimenting this, Wang et al. (Wang et al., 2023) demonstrated that using similar examples and increasing the number to four enhances performance, while adding more than four examples yields only marginal gains. Their work focused on classification, multiple-choice, and generation tasks, rather than user review analysis tasks. Prakash et al. (Prakash et al., 2023) evaluated prompts with 2, 4, 6, and 8 examples, finding comparable topic coherence and diversity across configurations.
In this paper, we further extend this line of research by investigating how the number of in-context examples and the exclusion of noisy data affect topic extraction and assignment tasks for user review. Our work demonstrates that increasing the number of examples for topic extraction from 0-shot to 5-shot significantly improves the model’s accuracy. Our work also suggests that filtering out non-informative reviews enhances the accuracy.
3. Methodology
To study user feedback on Gen-AI apps, we design a four-phase methodology we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning pipeline to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. Our proposed methodology enables us to address our RQs concerning the impact of data quality and few-shot examples on LLM performance, the prevalent topics within Gen-AI user reviews, and the temporal evolution of user ratings toward these apps. We present an overview of our research approach in Figure 2.
Figure 2. Overview of our approach
3.1. Selection of Gen-AI apps
Our app selection process involves three steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, and 3) manually screening and validating each app to confirm it is Gen-AI.
  1. 1.
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Garrido-Merchán, 2023). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
  2. 2.
    App scraping. Using our list of keywords, we scrape the Google Play Store to identify all matching apps. Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store. For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description. In this paper, we refer to Google Play Store app categories as app categories.
  3. 3.
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
Table 1. Categories and keywords for app search in the Google Play Store
Category Keywords
Text Conversational AI, Text-to-Science, Text-to-Author Simulation, Text-to-Medical Advice, Text-to-Itinerary, Doc-to-Text
Image Image Editing, Artistic Images, Realistic Images
Video Text-to-Video, Video Production, Image-to-Video
3D Text-to-3D
Code and Software Text-to-Code, Text-to-Multilingual Code
Speech Text-to-Speech, AI Understanding, Speech-to-Text
Business Business
Marketing Marketing, Advertisement, Accounting
Gaming Videogame Creation
Music Music Generator
3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, nd). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning pipeline designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
Table 2. Number of reviews by app category and stage.
App category Stage 0 Stage 1 Stage 2 Stage 3
Productivity 362,210 354,829 233,311 169,119
Photography 1,035,946 430,490 188,382 126,277
Art & Design 332,881 238,291 157,320 125,362
Entertainment 205,467 205,467 156,636 118,669
Education 101,812 95,590 57,680 41,905
Video Players & Editors 364,124 109,610 51,182 36,507
Tools 231,865 104,636 48,807 35,046
Music & Audio 17,875 17,875 11,781 9,150
Health & Fitness 41,537 10,368 8,747 7,849
Games 29,922 6,254 5,555 5,070
Books & Reference 2,071 2,071 1,423 1,112
Total 2,725,710 1,575,481 920,824 676,066
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labelled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar et al., 2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi and Elluri, 2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar, 2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021a; Arambepola et al., 2024). This step aims to eliminate overly brief and potentially non-informative feedback.
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
  1. 1.
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024b).
    We design a filtering prompt that instructs the LLM to classify reviews as either informative or non-informative. To guide the model effectively, we incorporate representative examples within the prompt.
  2. 2.
    Iterative prompt refinement. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labelled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.
    Figure 3. Template of the filtering prompt used to filter out non-informative reviews.
  3. 3.
    Prompt validation. To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size. The model demonstrates strong generalization capabilities, achieving an average classification accuracy of 90%.
  4. 4.
    Large-scale review filtering. We apply the validated filtering prompt to the full dataset to filter all non-informative reviews. Reviews are processed in batches of 100 to minimize omissions, as larger batches sometimes result in the model returning only a subset of reviews. Any unreturned reviews are consolidated into new batches and resent. This process reduces 920,824 reviews to 676,066 informative reviews, as shown in Table 2.
Our data collection, spanning September and October 2024, resulted in an initial dataset of 2,778,643 user reviews across 186 unique Gen-AI apps. Following our multi-stage cleaning process, we obtain a dataset of 676,066 informative reviews across 173 Gen-AI apps, which forms the basis for our subsequent analysis.
3.4. LLM-powered user review analysis
We use the refined user review data to conduct topic extraction, topic assignment, trend analysis, and manual analysis. For RQ1, we design and evaluate our topic extraction and assignment prompts. We experiment with different data noise levels (i.e., different stages) and numbers of few-shot examples. For RQ2, we extract the top 10 topics per app category and assign them to user reviews. We conduct manual thematic analysis for each Gen-AI topic to derive user insights from reviews. For RQ3, we analyze how Gen-AI topics evolve over time using clustering and manual review. In the following sections, we describe the experiment design and the findings for each RQ.
4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin et al., 2024). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning pipeline. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar, 2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
Figure 4. An overview of our RQ1 experiment design, where $x\in\{Stage1,Stage2,Stage3\}$ and $y\in\{Photography,Productivity,Art\&Design,Entertainment,VideoPlayers\&Editors\}$.
Sampling. To evaluate the effectiveness of our topic extraction prompts across different stages of data cleaning, we define $x\in\{Stage1,Stage2,Stage3\}$ to represent the three data cleaning stages. For each stage $x$, we select five representative app categories from the Google Play Store, i.e., Photography, Productivity, Art & Design, Entertainment, and Video Players & Editors, denoted as $y$. For each combination of cleaning stage $x$ and app category $y$, we generate two separate statistically representative random samples of user reviews from our review dataset. First, we extract a large sample denoted $S_{\text{large}}(x,y)$, computed using a 95% confidence level and a 2% margin of error with an average size of 2,337 reviews. This large sample is used for topic extraction to ensure the coverage of the user feedback space. Second, we draw a small sample denoted $S_{\text{small}}(x,y)$ from our review dataset, computed using a 95% confidence level and a 10% margin of error with an average size of 96 reviews. This sample is used for topic assignment, where manual evaluation is feasible and interpretable.
Topic extraction. To evaluate how prompt design affects topic extraction performance, we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instruction with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
Figure 5. Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$. Insert category name is where we insert the name of the app category $y$. Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the set of top topics $T(x,y,n)$ extracted in that experiment, 2) five few-shot example reviews manually selected and labelled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025).
Figure 6. Template of the P5: topic assignment prompt used to assign topics to a list of reviews. We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labelling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000). We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
4.2. RQ1 Findings
Observation 1.1: Providing more examples significantly improves classification and assignment accuracy. The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3. Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples. The Kruskal–Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$). Post hoc Mann–Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not. These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
Table 3. Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 86% 77% 91% 69% 95% 83% 84% 85% 85% 80% 90% 85% 72% 86% 77% 86% 89% 82%
E2: 3 shot 91% 86% 80% 66% 95% 83% 91% 80% 86% 82% 95% 87% 76% 83% 88% 92% 90% 86%
E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy. As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3). For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3. In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3. This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3. Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy. As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments. Given the inconsistencies observed in overall accuracy explained in observation 1.2, we isolate informative reviews to better evaluate model performance. Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3. This suggests that data cleaning improves the model’s ability to assign topics to reviews with informative content correctly. We conduct a one-way ANOVA test to evaluate whether the cleaning stage has a significant effect on classification accuracy. The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 83% 76% 90% 88% 76% 83% 84% 83% 89% 74% 82% 82% 81% 91% 81% 86% 89% 86%
E2: 3 shot 87% 91% 87% 86% 76% 85% 93% 78% 89% 89% 94% 88% 81% 88% 89% 91% 90% 88%
E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot. Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model. Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
Summary of RQ 1 Our findings indicate that increasing the number of few-shot examples in the prompt significantly improves output accuracy. Additionally, filtering out non-informative reviews enhances the accuracy of extraction and assignment for informative reviews. The model using the 5-shot prompt combined with LLM-based filtering of non-informative reviews achieves the highest overall accuracy of 91%. Accordingly, we adopt this configuration for topic extraction in RQ2.
5. RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
5.1. Experiment Setup
In RQ1, we experiment with varying levels of data cleaning and numbers of few-shot examples. We find that filtering out non-informative reviews using an LLM and providing five-shot examples improves topic extraction and assignment accuracy. In RQ2, we apply the optimal topic extraction prompt identified in RQ1 to the remaining app categories and assign topics to all reviews in Stage 3. In this section, we provide a detailed description of the implementation of RQ2.
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totalling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform quantitative analysis on both app categories and topic categories. We compute the following metrics:
  1. (1)
    Average Gen-AI Rating (AGR): the average star rating of reviews labelled with Gen-AI topics.
  2. (2)
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labelled with non-Gen-AI topics.
  3. (3)
    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI
  4. (4)
    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics
  5. (5)
    Gen-AI Review Percentage (GRP): the proportion of each type relative to the total number of reviews in each app category
  6. (6)
    Non-Gen-AI Review Percentage (NRP): the proportion of each type relative to the total number of reviews in each topic category
We conduct the Mann–Whitney U test to examine whether reviews labelled as Gen-AI-related receive significantly higher ratings than those that are not.
To gain deeper insights into user concerns and feedback within each topic category, we randomly sample reviews using a 95% confidence level and a 10% margin of error, resulting in a dataset of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality exceeds the average rating of non-Gen-AI reviews in every app category, indicating higher user satisfaction with the Gen-AI features.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups ($U=65{,}113{,}513{,}850$, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
Table 5. Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics
App Category AGR ANR GRC GRP # Gen-AI Topics
Productivity 4.2 3.5 84,122 50% 6
Entertainment 4.0 3.2 56,672 48% 6
Education 4.6 3.3 31,465 75% 6
Art & Design 3.4 3.2 26,585 21% 3
Photography 3.9 3.3 15,567 12% 2
Tools 4.0 3.4 7,782 22% 3
Health & Fitness 4.8 4.0 6,896 88% 7
Video Players & Editors 4.6 3.9 5,179 14% 2
Games 3.6 2.7 1,788 35% 5
Music & Audio 3.0 2.6 1,679 18% 4
Books & Reference 4.1 2.3 428 39% 4
Average 4.0 3.2 21,651 38% 4
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As reported in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. For example, categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively. Conversely, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews. In these categories, 58% of the reviews focus on non-Gen-AI topics, such as overall user experience, technical issues, and subscription or pricing models, suggesting that these factors outweighed interest in Gen-AI features.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting key differences in engagement. Several categories, such as Education and Productivity, exhibit a moderate spread, indicating consistent Gen-AI engagement across apps. Other categories, such as Art & Design and Entertainment present a wider spread, suggesting more variability in Gen-AI adoption among apps within these groups.
Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial tool for boosting overall user satisfaction. They must strategically incorporate Gen-AI features to remain competitive.
Observation 2.3: The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. In terms of user satisfaction, the Gen-AI topics with the highest average ratings are Human AI Interaction & Emotional Connection, Utility & Use Cases, and Customer Support & Community. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. We further examine each of these categories in the subsections below.
Table 6. Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears
Gen-AI Topic Category GRC AGR # Apps # App Categories
AI Performance 61,071 4.0 127 10
Utility & Use Cases 56,505 4.8 45 3
Content Quality 44,594 3.6 131 8
Content Policy & Censorship 27,521 2.9 76 5
Human AI Interaction & Emotional Connection 25,142 4.8 13 4
Features & Functionality 16,275 4.0 58 4
Customer Support & Community 3,361 4.7 12 3
Comparison to Other Apps 2,315 3.9 13 3
Personalization & Customization 1,283 4.0 30 3
Accessibility & Inclusivity 96 4.1 4 1
Table 7. Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears
Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
Technical Difficulties 99,621 2.2 157 10
Monetization Methods & Structure 85,756 2.3 159 11
User Interface & Experience 80,827 4.5 156 11
Utility & Use Cases 21,464 4.8 34 1
Updates & Evolution 21,100 2.8 92 7
Personalization & Customization 6,860 4.8 18 1
Customer Support & Community 6,848 1.8 85 6
Limitations of Usage 5,971 2.5 45 4
Features & Functionality 1,810 3.0 11 1
Comparison to Other Apps 1,493 3.4 17 1
Accessibility & Inclusivity 401 2.6 5 2
Other (unassigned) 105,752 4.2 165 11
1) AI Performance. AI Performance refers to how well the AI model performs in terms of 1) accuracy (27%), 2) speed (22%), 3) understanding (15%), 4) memory (5%), 5) reliability (5%), and 6) creativity & randomness (3%). AI accuracy describes the precision of the AI response to the user’s prompt. AI speed represents how quickly the AI responds to a user’s prompt. AI understanding indicates the AI’s ability to comprehend the meaning, intent, and context behind the user’s prompt. AI memory signifies how well the AI remembers prior conversations and links them to the current one. AI can have both short-term memory, which spans a single conversation, and long-term memory, which spans multiple conversations. Reliability refers to the consistency and dependability of the AI system in performing its intended tasks without crashing, freezing, timing out, or failing to generate output. Creativity captures the degree of originality, uniqueness, or unpredictability in the AI’s responses.
Figure 8. Examples of reviews discussing AI Performance
Insights. We find that 79% of the users who discussed AI Performance provided a positive review with a rating of 4 or 5. Reviews 1 and 2 in Figure 8 show examples of positive reviews. Despite this overall satisfaction, 16% of reviews mentioned limitations in AI Performance, with the most frequently cited issue being AI understanding (5%), followed by inconsistency in accuracy (4%), reliability (3%), speed (2%), and AI memory (1%).
A recurring theme in the less favourable reviews is poor AI understanding, which leads to irrelevant or partially correct responses. Some users show awareness of prompt engineering and attempt to guide the AI with examples, yet still encounter issues. For instance, review 3 in Figure 8 illustrates a user struggling to communicate effectively with the AI, despite providing clear examples. Similarly, review 4 in Figure 8 highlights inconsistent performance for complex requests. While the AI excels at handling simple tasks, it often struggles with more complex ones.
Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
$\Rightarrow$ Implication 2.
While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
Figure 9. Examples of reviews discussing Content Quality
Insights. In our sample, 67% of reviews express satisfaction, praising the creative and polished outputs such as producing “radio-ready” songs and appending useful resources to textual answers as illustrated in reviews 1 and 2 in Figure 9 respectively.
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk et al., 2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 3.
Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
3) Content Policy & Censorship. Content policy refers to rules and guidelines that define what type of AI-generated content is permissible. Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those rules.
Figure 10. Examples of reviews discussing Content Policy & Censorship
Insights. In our review sample, 73% of users comment on censorship experience. Of these 23% report frustration with the filtering that often block benign content, adversely affecting usability. For example, the author of review 1 in Figure 10 reports getting censored for saying “hi.” On the other hand, 10% of the raised concerns relate to lack of effective filtering such as unethical imagery involving children bypassing filters as expressed in review 2 in Figure 10. Regarding adult content, 67% of concerned users express dissatisfaction with the excessive restriction of access. While 19% explicitly request an optional censorship toggle for adult users as illustrated in review 3 in Figure 10. Additionally, users express concern over unclear content ownership. Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
$\Rightarrow$ Implication 4.
Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content. Ideally, developers could offer optional and adjustable censorship settings for mature or context-specific use. At the same time, platforms must clearly state who owns AI-generated outputs and how they can be reused, to maintain transparency and protect user trust.
4) Features & Functionality. Features & Functionality refer to the Gen-AI-driven capabilities offered by the app, such as voice interaction and prompt autocompletion, which shape the user experience and determine which features users find valuable or lacking.
Insights. In our sample, 75% of user reviews praise specific Gen-AI features, while 25% criticize them. For example, some users react negatively when a “stop-generation” button is removed and replaced with a voice mode feature, as illustrated in review 1 in Figure 11. While 11% of users enjoy the voice features, some describe them as useless. This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
$\Rightarrow$ Implication 5.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
Figure 11. Examples of reviews discussing Features & Functionality
5) Utility & Use Cases. Utility & Use Cases refer to the practical roles Gen-AI plays in users’ daily lives, including enhancing accessibility, supporting professional tasks, and enabling creative expression across diverse contexts.
Figure 12. Examples of reviews discussing Utility & Use Cases
Insights. In our sample, 64% of reviews emphasize AI’s educational utility, such as helping with homework, essay writing, tutoring, and breaking down complex topics as shown in review 1 in Figure 12. Additionally, some users specifically praise how Gen-AI can act as an accessible on-demand learning assistant for dyslexic users, as commented in review 2 in Figure 12. Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
$\Rightarrow$ Implication 6.
Despite the launch of tools like ChatGPT Edu (OpenAI, 2024a), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
6) Personalization & Customization. Personalization & Customization describe how Gen-AI systems adapt to user preferences and allow users to tailor outputs. This includes customizing output formats, editing generated content, and adopting preferred AI personas, e.g., voices or AI character roles.
Figure 13. Examples of reviews discussing Personalization & Customization
Insights. In our sample, 38% of reviews emphasize the AI’s inability to create a personalized experience. For example, review 1 in Figure 13, discusses the app’s inability to maintain shopping and to-do lists. Additionally, 10% of users request inline editing capabilities to support customization by refining generated content directly, rather than having to regenerate entirely. For instance, review 2 in Figure 13 suggests the option to edit lyrics in an already generated song rather than regenerate a new one entirely. Meanwhile, 9% of reviews request more diversity in AI characters and custom voices. This reflects a growing interest in personalizing AI identities to match preferences such as female voice options. For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
$\Rightarrow$ Implication 7.
Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized.
7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
Figure 14. Examples of reviews discussing Comparisons to Other Apps
Insights. We find that 66% of the reviews draw comparisons to other AI-powered apps (e.g., comparing Microsoft Bing Search with Google). We also observe that 9% of the reviews compare the app with non-Gen-AI alternatives. For instance, review 1 in Figure 14 criticizes the forced integration of the AI chat feature, expressing a strong preference for traditional search functionality. Finally, 18% of reviews draw comparisons between the apps they review and general-purpose AI tools (e.g., ChatGPT), highlighting expectations shaped by user experiences with AI tools, as shown in review 2 in Figure 14. 77% of the reviews that compare the app with ChatGPT believe that ChatGPT outperforms the app in question, while 15% prefer the performance of the domain-specific app. These comparisons highlight a key challenge for domain-specific Gen-AI app developers. Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
$\Rightarrow$ Implication 8.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).
8) Customer Support & Community. Customer Support & Community refers to reviews that address developers directly to express their experience, such as reporting user enjoyment with specific features or requesting to fix app shortcomings.
Figure 15. Examples of reviews discussing Customer Support & Community
Insights. In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
$\Rightarrow$ Implication 9.
User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Assi et al., 2025a; Hassan et al., 2022; Assi et al., 2021b). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
9) Human AI Interaction & Emotional Connection. Human AI Interaction & Emotional Connection reviews highlight the emotional and relational dimensions of human–AI interaction, particularly with conversational Gen-AI apps.
Figure 16. Examples of reviews discussing Human AI Interaction & Emotional Connection
Insights. We observe that users enjoy their interactions with conversational Gen-AI apps for multiple reasons. First, users enjoy a realistic interaction (i.e., closely similar to human dialogue). This perceived realism fosters a sense of companionship and emotional connection (e.g., review 1 in Figure 16). Second, reviews highlight the comfort in the emotional support provided by Gen-AI apps, particularly because the interaction lacks real human judgment, as shown in review 2 in Figure 16. This illustrates the importance of AI’s perceived neutrality and non-judgmental nature, especially when discussing sensitive or emotionally charged topics. Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
$\Rightarrow$ Implication 10.
The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability. Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care. Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
10) Accessibility & Inclusivity. Accessibility and inclusivity refer to the degree to which the Gen-AI apps accommodate users for affordability, multilingual support, age appropriateness, and usability for individuals with cognitive or physical impairments.
Figure 17. Examples of reviews discussing Accessibility & Inclusivity
Insights. Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
$\Rightarrow$ Implication 11.
These findings highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents. They expect fast, creative, and context-aware responses, along with personalization and transparency. For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT).
6. RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
To understand how users perceive Gen-AI features over time, we explore the temporal dynamics of user feedback and ratings in app reviews. Our goal is to uncover how user opinions about Gen-AI topics shift as Gen-AI capabilities mature and become more integrated into mobile apps.
6.1. Experiment Setup
We track the evolution of user feedback over time by conducting a two-level trend analysis. First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time. This perspective helps highlight the increasing user attention to Gen-AI features. Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns. We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis. Based on the observed rating dynamics, we select three representative topic categories for deeper investigation: Content Quality (decreasing trend), AI Performance (stable trend), and Content Policy & Censorship (increasing trend). The following steps outline how we conduct the analysis:
Step 1: data preparation. We group reviews according to their topic category $(T)$ and time period $(P)$. Specifically, for every topic category $T$, we organize reviews into half-year intervals (i.e., every 6 months) based on their submission dates. Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of reviews in period $P$ that belong to category $T$. We exclude the topic category Accessibility & Inclusivity because it only contains 96 reviews in total, which are not enough to conduct a time series analysis.
Step 2: clustering methodology. To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends. For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we extract slope and standard deviation as features following the three-step process listed below:
  1. 1.
    Features extraction. Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash and Koot, 2009). Feature selection enhances interpretability, aligns with research goals  (Kreienkamp et al., 2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable). For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$. Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend. It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
  2. 2.
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren Liao, 2005). We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b). K-Means is suitable for low-dimensional feature spaces (Ikotun et al., 2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
  3. 3.
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari et al., 2023), across multiple values of $k$. Based on the silhouette results, we select the $k$ value closest to 1. The optimal silhouette scores correspond to $k$=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$. We complement this with qualitative validation through manual inspection, confirming that $k$=3 yields the most interpretable and distinct groupings.
Step 3: manual trend investigation. To interpret the identified clusters, we perform qualitative analysis on random review samples from key time periods. For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline. From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error. We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations. This process provides qualitative context that enhances our understanding of the observed quantitative trends.
6.2. Findings
Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time. Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics. As illustrated in Figure 18, the percentage of reviews discussing Gen-AI topics increases over time. The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features. Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 19 with only a slight change of ± 0.2 stars across the observed period. Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories. As illustrated in Figure 20, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable. These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings.
Figure 18. Evolution of the Percentage of Gen-AI Reviews
Figure 19. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics
Gen-AI topics analysis. As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters. These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 20 and 21 for the average ratings and the percentage of reviews, respectively.
To deepen our understanding of the observed rating trends, we perform a manual trend investigation, i.e., a qualitative analysis, on three representative topic categories, each corresponding to a distinct trend direction. For the Decreasing trend, we select Content Quality due to the unexpected nature of its decline, particularly in light of ongoing advancements in AI capabilities as shown in Figure 20. For the Stable trend, we examine AI Performance, where we had anticipated an upward trend reflecting improvements in system performance. For the Increasing trend, we analyze Content Policy & Censorship to investigate how user perceptions of policy changes have evolved and to uncover the factors contributing to rising satisfaction in this area.
Figure 20. Trend Clusters for Average Ratings of Gen-AI Topics
Figure 21. Trend Clusters for Percentage of Gen-AI Topics
Observation 3.2: The Content Quality topic category shows a decline in rating over time, despite improvements in Gen-AI technology. As shown in Figure 20, the average rating for Content Quality decreases from 4.5 in 2022-H1 to 3.3 in 2024-H2. To better understand the unexpected downward trend in Content Quality ratings, we conduct a manual analysis of user reviews across three strategically selected time periods: 2022-H1 (corresponding to the highest average rating), 2022-H2 (marking a period of sharp decline), and 2024-H2 (representing the lowest observed average rating). This sampling allows us to investigate user behavior before, during, and after the major shifts in rating averages. Based on the manual inspection, we identify two key user-related factors driving the decrease as follows.
Figure 22. Examples of reviews discussing Content Quality at different time periods.
Observation 3.3: Over time, the Content Quality reviews become less tolerant of flaws and more likely to penalize them. Excitement for emerging Gen-AI technologies appears to influence early adopters’ evaluation criteria, making them less likely to penalize imperfections. In our analysis of the 2022-H1 sample, 96% of the reviews are positive, with only 5% of them including constructive feedback despite assigning a perfect 5-star rating. For example, the user writing review 1 as shown in Figure 22 awards the app the highest rating while reporting that the AI generated low-quality facial images. Such reviews reflect a pattern in which users prioritize innovation and novelty over technical quality, demonstrating a higher tolerance for flaws during the early adoption phase.
Over time, users become less tolerant of imperfections. The percentage of positive reviews declines from 96% in 2022-H1 to 59% in 2024-H2, indicating a significant shift in user ratings. Moreover, positive reviews become increasingly critical: in 2022-H2, 12% of all reviews offer constructive feedback while still assigning a high rating (4 or 5 stars), rising to 18% in 2024-H2. This trend suggests that users’ expectations have heightened.
More notably, an increasing number of users begin deducting rating star points for flaws over time. In 2022-H2, only 2% of constructive reviews assign a 5-star rating, while 10% assign a 4-star rating. By 2024-H2, this trend intensifies: 15% of constructive reviewers assign a 4-star rating, whereas only 4% assign a 5-star rating. For instance, review 2 in Figure 22 expresses the need for improvement in graphics and gives a 4-star rating. This shift presents a growing criticality among users. While early reviews tend to overlook shortcomings in favour of innovation, later reviews are more discerning, penalizing points even when expressing generally positive sentiment.
Observation 3.4: Over time, the Content Quality reviews are more experimental and demanding. Initial reviews frequently convey surprise and satisfaction with relatively modest capabilities, indicating limited expectations during the early adoption phase. For example, review 3 in Figure 22 from 2022-H1 expresses surprise at the high quality of the content, despite initially having low expectations. Such statements suggest that early positive ratings are often driven by low expectations of the novel technology rather than its actual performance.
As Gen-AI technologies become more widespread, user expectations rise accordingly. Reviews begin to reflect more ambitious and exploratory usage patterns, with users increasingly testing the system’s capabilities through complex prompts. When the app fails to meet these heightened demands, users express dissatisfaction. For example, review 4 in Figure 22 in 2022-H2, reports success in simple requests but horrific results with complex requests, ultimately giving the app a poor 1-star rating. This evolution in review content presents a shift from initial awe to critical evaluation, as users become more experimental and less tolerant of limitations.
We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar, 2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity. Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino and Gudmundsson, 2024). Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
$\Rightarrow$ Implication 12.
Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users. Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1. To account for this overall stability, we analyze fluctuations at the app category level. We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others. This inter-category balancing effect contributes to the relatively flat trend observed in the aggregate ratings.
Figure 23. Examples of reviews discussing AI Performance at different time periods.
Observation 3.6: Advances in image generation capabilities contribute to the slight increase in the average rating of the AI Performance topic in 2023-H1. In 2022-H1, negative reviews of image generation apps constitute 9% of reviews in our sample and account for 47% of all negative reviews during that period. These users express frustration with the visual output. For example, review 1 in Figure 23 expresses frustration over the performance of the image generator.
By 2023-H1, these issues appear to be addressed, as the percentage of negative reviews of image generation apps drops to 5%, and their share of total negative reviews decreases to 29%. This reduction in dissatisfaction likely suggests improvements in the AI Performance of image generation, which helps lift the overall AI Performance rating during this period.
Observation 3.7: Underperformance in Productivity apps leads to a slight decline in the average rating of the AI Performance topic in 2024. By 2024-H1, Productivity apps emerge as the leading source of performance-related complaints, accounting for 8% of all reviews in our sample and 44% of all negative reviews. Unlike creative tools, which are often evaluated in terms of novelty or expressiveness, Productivity apps are frequently judged against well-established, non-AI alternatives that offer straightforward functionality. Consequently, users express notable frustration with AI-driven features that are perceived as unnecessarily complex or unreliable. For example, review 2 in Figure 23 prefers a regular assistant over an AI assistant because the AI assistant complicates simple tasks, and review 3 in Figure 23 prefers a calculator over an AI due to accuracy issues.
These comments highlight that Gen-AI integration can complicate tasks that would otherwise be simple. Rather than enhancing the user experience, AI features may undermine usability if they are poorly aligned with user expectations or fail to deliver practical value. Some reviews explicitly question the necessity of AI features. Review 2 in Figure 23 describes abandoning their AI assistant after it offered a detailed classification of fan technologies instead of simply turning on a fan. This indicates that forced integration may backfire, especially when it disrupts core functionality.
Observation 3.8: The stability in the average rating of the AI Performance topic is driven by offsetting trends across app types. Some app categories show improvement (e.g., image generation), while others show decline (e.g., Productivity). These shifts tend to cancel each other out at the global level. In 2023, rising satisfaction with creative apps contributes to a mild increase in performance ratings. In 2024, dissatisfaction in Productivity apps is balanced by relatively stable perceptions in creative and educational apps. This compensatory effect explains the flat trend in average performance ratings over time.
Although the global average for AI Performance ratings remains stable over time, this apparent stability masks meaningful variation across app categories. Our analysis reveals divergent trends: categories such as Art & Design, Entertainment, and Books & Reference show moderate declines in average AI Performance ratings (–0.84 to –0.93), while others like Education and Games experience notable increases (+1.39 and +0.96, respectively). These opposing shifts suggest that stability at the aggregate level may result from the balancing of improvements and declines across distinct domains. This trend may also reflect differences in user expectations. Creative tools are often evaluated for imagination and output quality, while Productivity apps are judged more on efficiency and reliability.
$\Rightarrow$ Implication 13.
AI integration should be purpose-driven and context-aware. Raji et al. (Raji et al., 2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility. The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration. Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default. Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1. We observe that user ratings decrease in 2022-H2, rise steadily through 2023-H2, and decrease slightly in 2024-H1. To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1. We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
Figure 24. Examples of reviews discussing Content Policy at different time periods.
Observation 3.10: Ethical concerns around AI-generated art are more prominent in early samples but disappear in later ones. In 2022-H2, 10% of sampled reviews express concerns about the ethics of training AI on copyrighted human art without consent or compensation. For example, review 1 in Figure 24 expresses how Gen-AI harms artists whose art is stolen. This concern disappears from the random samples of 2023-H2 and 2024-H1.
This highlights the importance of transparent and ethical data use policies. However, the absence of similar comments in later samples suggests either that the user-base shifted or that these concerns were deprioritized in users’ feedback. The drop may also indicate user desensitization or limited platform changes aimed at addressing the issue. Nevertheless, the ethics of AI content policy should still be addressed by researchers and developers.
Observation 3.11: Content filters improve over time in blocking unwanted adult content (i.e., the raised concerns drop from 12% to 3%) and illegal content (i.e., the raised concerns drop from 4% to 0%). In our early sample, 15% of the studied reviews flag harmful content slipping past moderation, especially involving sexual or predatory material. This drops to 5% in 2023-H2 and to 3% by 2024-H1, indicating increased user satisfaction with content safety mechanisms. Examples of such reviews can be found in reviews 2, 3, and 4 in Figure 24.
The proportion of users reporting the unintentional generation of adult content drops from 12% in 2022-H2 to 3% in 2024-H1. Reports of illegal or predatory content involving children drops from 4% to 0%. These declines reflect measurable improvement in censorship mechanisms for harmful content.
Observation 3.12: Complaints about false-positive censorship grow over time (from 35% to 59%), especially among users engaging in regular and appropriate use. As filters improve at blocking harmful content with time, users increasingly report unnecessary censorship during benign interactions. In 2022-H2, 35% of users in our sample request a less strict filter. This jumps to 71% in 2023-H2 and settles at 59% in 2024-H1.
While the proportion of users requesting more freedom to generate explicit or violent content decreases slightly (from 21% in 2022-H2 to 14% in 2024-H1), more users begin to complain about filters interfering with regular usage. For example, 6% of reviews in 2022-H2 flag such issues, rising to 30% in 2024-H1. Review 5 in Figure 24 shows an example of reviews discussing false-positive censorship, where the user reports receiving a guidelines warning for writing a non-explicit role-play scene. This suggests that while filters are more successful in removing harmful content, they are over-correcting and obstructing benign, creative, or functional uses. Over-correction can hinder the Gen-AI user experience, particularly in role-play and storytelling apps.
Observation 3.13: Despite an overall positive trend, over-correction in censorship filters causes a minor ratings setback in 2024. The average rating for Content Policy and Censorship follows an increasing trend over time, reflecting growing user trust in content moderation systems. As previously discussed, this trend is supported by a substantial decline in complaints about harmful content. Correspondingly, the share of positive reviews that discuss censorship rises from 4% in 2022-H2 to 51% in 2023-H2. This rise indicates that while many users enjoy the app enough to leave a high rating, they still express frustration with the limitations imposed by the censorship filters.
However, this progress introduces a new challenge of false-positive filtering in 2024-H1. As previously discussed, while moderation systems are successful in blocking inappropriate content, they begin flagging benign prompts as violations. This over-correction results in a moderate decline in the share of positive censorship-related reviews (20% in 2024-H1) and contributes to a slight drop in the average rating. Nevertheless, this dip does not reverse the overall gains. Ratings in 2024-H1 remain higher than in 2022-H2, suggesting that while usability concerns emerge, users still view the improvements in content safety as a net positive. This also highlights an important reality for developers; progress in content moderation is unlikely to be perfectly linear. Minor setbacks and fluctuations are to be expected as systems evolve and adapt to user needs.
$\Rightarrow$ Implication 14.
Developers must strike a careful balance between safety and creative freedom. While stricter filters reduce the occurrence of disturbing or illegal content, they increasingly disrupt legitimate use. To avoid over-correction, developers should implement modular moderation systems with customizable filters or toggles that can be adjusted by personal preference. Clear content policy guidelines and transparent communication about AI behaviour are also essential. Additionally, ethical considerations around AI training data should not be ignored by developers or researchers, even if fewer users express concerns. These considerations remain essential for maintaining long-term trust and ensuring platform responsibility.
Summary of RQ 3 Our analysis reveals that Gen-AI topics evolve differently over time with increasing, stable, or decreasing trends. Content Quality ratings steadily decline, not due to technological regression, but due to increased user expectations. In contrast, AI Performance shows a stable rating trend, driven by a balancing effect between improving and declining domains. Lastly, Content Policy & Censorship follows an overall positive trend, with improved content filters increasing user ratings. These findings reflect a maturing user base that is more experienced, discerning, and vocal. Our user base of Gen-AI apps raise the bar for what constitutes a high-quality Gen-AI experience.
7. Discussion
In this section, we reflect on the broader implications of our findings and provide targeted recommendations for six key stakeholder groups: 1) users, 2) platform owners, 3) developers, 4) testing tool designers, 5) requirements engineers, and 6) policymakers and regulators.
7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 2, 7, and 10). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 21).
7.2. Implications for Platform Owners
Platform owners must bear responsibility for ensuring that Gen-AI apps comply with ethical standards (Observation 3.10). As AI capabilities evolve rapidly, app store operators should implement both technical tools and regulatory frameworks that evaluate whether submitted Gen-AI apps meet baseline criteria for safety, transparency, and responsible data usage. This may include requiring apps to disclose whether AI is used, clarifying how outputs are generated, and verifying that training data sources are ethically obtained. By introducing certification processes or audit mechanisms for Gen-AI features (particularly in apps that claim educational, therapeutic, or professional utility), platforms can promote accountability and mitigate harm, fostering greater trust among users and developers alike.
App platforms play a critical role in shaping the quality of user feedback. Our results show that a large proportion of user reviews are non-informative (Observation 1.2), which limits their value for developers and researchers. Future research could explore how redesigning review interfaces might encourage more meaningful feedback. One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 8). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 6 and 10). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 4), supporting accessibility and inclusivity concerns (Implication 11), and enabling more personalized and flexible user experiences (Implication 7). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 13).
7.4. Implications for Testing Tool Designers
Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Observation 1.4), demonstrating the effectiveness of prompt design and data pre-filtering (Observations 1.1 and 1.3). These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows. Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment. Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles.
7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis approach provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality.
7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 4). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 4), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
8. Threats to Validity
Construct threats to validity refer to the risk that the measures we use in our study do not accurately reflect the concepts we aim to study. First, we use the average app review rating as a proxy for user satisfaction. However, this metric may not fully capture the sentiment expressed in the review text. For example, users sometimes assign high ratings while simultaneously reporting critical issues or limitations. To mitigate this threat, we supplement the quantitative rating data with manual qualitative analysis of review content, ensuring that nuanced sentiments, both positive and negative, are appropriately captured. Second, our methodology relies on LLMs to automatically filter and classify reviews. To control for potential biases introduced by the LLM-based automation, we manually annotate samples and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks. Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data. A key internal threat relates to the uncertainty surrounding the timeline for introducing the Gen-AI features app. While we manually verify that all selected apps currently support Gen-AI functionality, we cannot determine the exact date of integration. To reduce the risk of analyzing reviews unrelated to Gen-AI, we exclude all reviews prior to 2022, i.e., the year ChatGPT was released. While this filtering improves temporal alignment, it cannot fully guarantee that all remaining reviews pertain to post-Gen-AI functionality.
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,000 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
Our results show that LLMs can reliably extract user-discussed topics with high accuracy, i.e., 91%, when supported by five-shot prompting and data cleaning. We identify the 10 most frequently discussed topics revealing user behavior and challenges that inform actionable insights for developers and researchers. Furthermore, our findings suggest an evolution in user expectations, with a shift in user behavior, such as shifting from early adopters who were more forgiving of limitations to a mainstream user base that demands higher expectations. Future work will explore cross-platform comparisons and further investigate how expectations vary across different user profiles (e.g., casual vs. expert users) as Gen-AI tools continue and mature.
Data Availability. We provide a replication package, including manually labelled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.
Acknowledgements.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [RGPIN-2021-03969].