### v3

# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

Abstract.
The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps). Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice. In this work, we conduct a user-centered analysis of 676,066 automatically labeled reviews from 173 Gen-AI apps on the Google Play Store. We propose a structured four-phase framework, SARA (Selection, Acquisition, Refinement, and Analysis), which integrates and extends state-of-the-art techniques for large-scale review collection, filtering, and analysis using prompt-based LLMs. First, we empirically validate the reliability of LLMs for topic extraction and assignment, achieving 91% accuracy through five-shot prompting and LLM-based filtering of non-informative reviews. We then apply the framework to informative reviews to identify the ten most discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities. Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps. Based on our findings and observations, we present actionable implications for developers and researchers.
Generative AI, Large Language Models, User Review Analysis, Mobile Apps
$c$s: Software and its engineering Software maintenance tools
1. Introduction
Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod_Agal_2023). A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days. This rapid adoption signaled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces. In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat. Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel_Hartmann_Janiesch_Zschech_2024; 10628027). Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (gozalobrizuela2023surveygenerativeaiapplications).
Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi_Tashi_Qureshi_Shah_Muneer_Irfan_Zafar_Shaikh_Akhtar_Wu_et_al_2023). LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue. Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau_Hassan_Zhou_2025). When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen_Schlote_2024). A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs). APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen_Liu_Zhou_Zhao_Wang_Wang_Chen_Bissyandé_Klein_Li_2024).
While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges. Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen_Gao_Chen_Zhang_Liu_2025). These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao_Huang_Shen_Ma_Su_Wan_2024). Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern. Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024). As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
However, despite growing interest in integrating LLMs into mobile apps, little is known about the specific concerns and experiences of end-users. Existing research primarily focused on the challenges faced by developers. For example, Nahar et al. (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness. While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience. Chen et al. (Chen_Gao_Chen_Zhang_Liu_2025) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development. This study offers insight into implementation bottlenecks but also neglects user-facing concerns. Similarly, Shao et al. (Shao_Huang_Shen_Ma_Su_Wan_2024) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues. However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice. Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024; Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025; Yuen_Schlote_2024), creative image generation (Tang_Zhang_Ciancia_Wang_2024; haase2023art; Oppenlaender_Silvennoinen_Paananen_Visuri_2023), or mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025). Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas. To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains. We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration. This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
Figure 1. Examples of reviews of Gen-AI apps
To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps. For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration. Building on prior work in user review mining and LLM-based topic analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), we propose a structured four-phase framework called Selection, Acquisition, Refinement, and Analysis (SARA). This framework enables the systematic collection, cleaning, and analysis of user reviews from Gen-AI apps, demonstrating the feasibility of large-scale, reproducible studies using LLMs.
First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition). We then apply a data cleaning framework to enhance the quality of input data for LLM-based analysis (i.e., Refinement). Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis). Using this integrative framework, we automatically label and quantitatively analyze 676,066 user reviews from 173 Gen-AI apps using LLMs, complemented by qualitative analyses on statistically representative samples.
Before analyzing Gen-AI user reviews (RQ2–RQ3), we first evaluate the accuracy and reliability of our LLM-based framework (RQ1) to establish its validity for large-scale analysis. This ensures that the topic extraction and assignment process used in the subsequent analyses is accurate and trustworthy.
RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews. Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy. The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM framework for downstream analysis.
RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
To answer this RQ, we apply the LLM-based topic labeling framework from RQ1 to informative reviews across 173 Gen-AI apps. We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers. While our primary focus is on user reviews, we also include a small supplementary examination of developer responses to better contextualize how developers engage with Gen-AI feedback. Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset. Notably, some users form emotional bonds with conversational agents and value personalized interactions. These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
RQ3: What temporal trends emerge in user feedback on Gen-AI topics? In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures. We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation. Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time. These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
Our study provides valuable insights for app developers and researchers by uncovering how users perceive and evaluate Gen-AI features in mobile apps. Specifically, our contributions are as follows:
  • We propose a structured four-phase framework, SARA, that builds upon established techniques in user review analysis and LLM-based topic modeling. Using SARA, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve. While validated in the context of Gen-AI apps, SARA’s modular design makes it generalizable to analyzing user feedback in other mobile app domains and AI-integrated systems.
  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labeled examples and filtering out non-informative reviews significantly improves accuracy.
  • We provide actionable feedback for developers to maintain competitiveness, for policymakers to explore the ethical concerns surrounding content policy and censorship, and for user requirement engineers to better understand user needs.
  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labeled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
We organize the remainder of this paper as follows. Section 2 illustrates existing work on analyzing user reviews. Section 3 presents our methodology. Sections 46 describe the experiment design and the findings obtained for each RQ. Section 7 synthesizes our key findings and implications of our work. Section 8 discusses potential threats to the validity of our findings. Finally, Section 9 concludes the paper and discusses implications for future research.
2. Related Work
In our literature review, we examine two key areas: 1) the user perspectives of Gen-AI apps, and 2) user review analysis techniques.
2.1. User Perspective of Gen-AI Apps
The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024). Shata and Hartley (Shata_Hartley_2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption. Kim et al. (Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) compared student and faculty perceptions on the use of Gen-AI in learning. Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn. Both faculty and students believed that Gen-AI will have more negative than positive effects on learning. Yuen and Schlote (Yuen_Schlote_2024) surveyed adult learners to explore their views on the integration of AI in language learning apps. Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
In the creative image domain, Tang et al. (Tang_Zhang_Ciancia_Wang_2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools. They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs. In support of that, Haase et al. (haase2023art) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation. Oppenlaender et al. (Oppenlaender_Silvennoinen_Paananen_Visuri_2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey. Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (10842666; 10963316; 10985281; 10911418). Our focus is on Gen-AI-driven treatment through conversational agents. Jin et al. (Jin_Kim_Han_2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention. Heinz et al. (Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) studied the efficacy of a mental health AI chatbot. They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
Prior studies on user perceptions of Gen-AI apps have primarily relied on interviews and surveys to capture attitudes toward adoption, trust, and usefulness within specific domains such as education (Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) and image generation (Tang_Zhang_Ciancia_Wang_2024; Oppenlaender_Silvennoinen_Paananen_Visuri_2023). These works offer valuable insight into how users conceptualize Gen-AI, but they often reflect processed opinions rather than immediate experiences. In contrast, our analysis of large-scale user reviews captures lived interactions, representing users’ spontaneous reactions written in context and often within moments of use. This perspective complements survey-based research by revealing practical concerns and responses as they naturally occur. For example, while prior studies in mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) demonstrated that conversational AI can foster engagement and therapeutic rapport, our findings extend this understanding by showing that users explicitly value Gen-AI’s perceived neutrality, emotional support, and human-like realism. Furthermore, while earlier work examined single domains in isolation, our study provides a holistic view across multiple app categories, revealing shared topics and concerns that transcend context. By organizing insights by functional features rather than by domain, our findings highlight common usability, quality, and ethical challenges that characterize Gen-AI technologies more broadly.
2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis. These studies collectively form the foundation for our own framework SARA, which builds upon their methods but advances them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling. Ghosh et al. (Ghosh_Pargaonkar_Eisty_2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al_Wahshat_Abu_ulbeh_Yusoff_Zakaria_Amir_Fazamin_Wan_Hamzah_P_2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment. While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation. Our filtering component achieves comparable accuracy (90%) while operating at a much larger scale across diverse Gen-AI app categories.
Pre-LLM approaches to topic extraction. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi_Abran_2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (10499727) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modeling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi_Hassan_Tian_Zou_2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
These early techniques established the foundation for automated review analysis but relied heavily on statistical or lexicon-based models. With the advent of LLMs, topic extraction has shifted toward context-aware and semantically richer methods, as discussed next.
LLM-based topic extraction and assignment methods. LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola_Lalendra_Wimalasena_Munasinghe_2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham_Hoyle_Sun_Resnik_Iyyer_2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (openai2023gpt) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash_Wang_Hoang_Hee_Lee_2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews. Assi et al. (Maram_LLM_Cure) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Among these, LLM-Cure most closely resembles our framework in combining both extraction and assignment tasks, though its focus was on generating feature suggestions rather than large-scale trend analysis.
Single-task topic analysis. Instead of extracting and assigning topics to user reviews, some studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei_Courbis_Lambolais_Xu_Bernard_Dray_2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023) and Ren et al. (Ren_Nakagawa_Tsuchiya_2024) extracted topics from user reviews without performing the assignment task. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
Domain-specific applications of LLMs in user review analysis. Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis. Sorathiya and Ginde (Sorathiya_Ginde_2024) employed a hybrid approach combining Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss. Rezaei Nasab et al. (RezaeiNasab_Dashti_Shahin_Zahedi_Khalajzadeh_Arora_Liang_2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning based classifiers to detect fairness mentions and cluster them into six categories.
While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially. We analyze user reviews from a large-scale collection of Gen-AI apps spanning multiple functional categories, including but not limited to health & fitness, education, productivity, and entertainment. Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e. AI performance, content quality, and content policy) that are unique to this new class of applications. This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
Collectively, these studies demonstrate the growing potential of LLMs for review analysis yet reveal a gap in unified, scalable approaches that integrate filtering, topic extraction, and topic assignment across diverse app domains. To address this, we design the SARA framework, which operationalizes these steps into a cohesive process for large-scale analysis of Gen-AI user feedback.
3. Methodology
To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning framework to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. SARA integrates and extends state-of-the-art methods in user review analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), combining automated filtering, topic extraction, and topic assignment into a unified, empirically validated workflow that supports large-scale, LLM-driven analysis of user feedback. Although evaluated within the Gen-AI domain, SARA’s modular design enables its application to other mobile or AI-enabled systems for analyzing large-scale user feedback. We first evaluate the reliability of SARA in accurately labeling review topics (RQ1) and then apply it in subsequent analyses to explore user concerns and temporal trends (RQ2–RQ3). We present an overview of our research framework in Figure 2.
3.1. Selection of Gen-AI apps
Our app selection process involves three steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, and 3) manually screening and validating each app to confirm it is Gen-AI.
  1. 1.
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (gozalobrizuela2023surveygenerativeaiapplications). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
    Figure 2. Overview of our framework
  2. 2.
    App scraping. Using our list of keywords, we scrape the Google Play Store to identify all matching apps. Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store. For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description. In this paper, we refer to Google Play Store app categories as app categories.
  3. 3.
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (cohen1960kappa) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
Table 1. Categories and keywords for app search in the Google Play Store
Category Keywords
Text Conversational AI, Text-to-Science, Text-to-Author Simulation, Text-to-Medical Advice, Text-to-Itinerary, Doc-to-Text
Image Image Editing, Artistic Images, Realistic Images
Video Text-to-Video, Video Production, Image-to-Video
3D Text-to-3D
Code and Software Text-to-Code, Text-to-Multilingual Code
Speech Text-to-Speech, AI Understanding, Speech-to-Text
Business Business
Marketing Marketing, Advertisement, Accounting
Gaming Video game Creation
Music Music Generator
3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (googleplayscraper). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning framework designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
Table 2. Number of reviews by app category and stage (Full dataset considered).
App category Stage 0 Stage 1 Stage 2 Stage 3
Productivity 362,210 354,829 233,311 169,119
Photography 1,035,946 430,490 188,382 126,277
Art & Design 332,881 238,291 157,320 125,362
Entertainment 205,467 205,467 156,636 118,669
Education 101,812 95,590 57,680 41,905
Video Players & Editors 364,124 109,610 51,182 36,507
Tools 231,865 104,636 48,807 35,046
Music & Audio 17,875 17,875 11,781 9,150
Health & Fitness 41,537 10,368 8,747 7,849
Games 29,922 6,254 5,555 5,070
Books & Reference 2,071 2,071 1,423 1,112
Total 2,725,710 1,575,481 920,824 676,066
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh_Pargaonkar_Eisty_2024; Prakash_Wang_Hoang_Hee_Lee_2023; Roumeliotis_Tselikas_Nasiopoulos_2024; Lee_Lee_Seo_2024), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labeled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam_Ingio_2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar_Hasan_Kumar_Carroll_2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi_Elluri_2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar_2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi_Hassan_Tian_Zou_2021; 10499727). This step aims to eliminate overly brief and potentially non-informative feedback.
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen_Lin_Hoi_Xiao_Zhang_2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
  1. 1.
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (openai2023gpt) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI_O4Mini_2024).
    Prior research comparing explicit, implicit, and creative prompting (Giray_2023; Ho_Mayberry_Nguyen_Dhulipala_Pallipuram_2024; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning. Consistent with these findings and with approaches such as LLM-Cure (Maram_LLM_Cure), we design an explicit, structured prompt that classifies each review as either informative or non-informative. The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses. These measures ensure deterministic labeling and minimize hallucination by preventing free-form or extraneous text generation.
  2. 2.
    Iterative prompt refinement. Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; He_Rungta_Koleczek_Sekhon_Wang_Hasan_2024). In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (cohen1960kappa) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.
    Figure 3. Template of the filtering prompt used to filter out non-informative reviews.
  3. 3.
    Prompt validation. To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size. The model demonstrates strong generalization capabilities, achieving an average classification accuracy of 90%.
  4. 4.
    Large-scale review filtering. We apply the validated filtering prompt to the full dataset to filter all non-informative reviews. Reviews are processed in batches of 100 to minimize omissions, as larger batches sometimes result in the model returning only a subset of reviews. Any unreturned reviews are consolidated into new batches and resent. This process reduces 920,824 reviews to 676,066 informative reviews, as shown in Table 2.
Our data collection, spanning September and October 2024, resulted in an initial dataset of 2,778,643 user reviews across 186 unique Gen-AI apps. Following our multi-stage cleaning process, we obtain a dataset of 676,066 informative reviews across 173 Gen-AI apps, which forms the basis for our subsequent analysis.
3.4. LLM-powered user review analysis
We use the refined user review data to conduct topic extraction, topic assignment, trend analysis, and manual analysis. For RQ1, we design and evaluate our topic extraction and assignment prompts. We experiment with different data noise levels (i.e., different stages) and numbers of few-shot examples. For RQ2, we extract the top 10 topics per app category and assign them to user reviews. We conduct manual thematic analysis for each Gen-AI topic to derive user insights from reviews. For RQ3, we analyze how Gen-AI topics evolve over time using clustering and manual review. In the following sections, we describe the experiment design and the findings for each RQ.
4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; 10.1145/3736408). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In our setup, we adopt this principle by refining prompts through multiple pilot runs on statistically significant samples to improve label consistency and clarity. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning framework. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar_2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
Figure 4. An overview of our RQ1 experiment design, where $x\in\{Stage1,Stage2,Stage3\}$ and $y\in\{Photography,Productivity,Art\&Design,Entertainment,VideoPlayers\&Editors\}$.
Sampling. To evaluate the effectiveness of our topic extraction prompts across different stages of data cleaning, we define $x\in\{Stage1,Stage2,Stage3\}$ to represent the three data cleaning stages. For each stage $x$, we select five representative app categories from the Google Play Store, i.e., Photography, Productivity, Art & Design, Entertainment, and Video Players & Editors, denoted as $y$. For each combination of cleaning stage $x$ and app category $y$, we generate two separate statistically representative random samples of user reviews from our review dataset. First, we extract a large sample denoted $S_{\text{large}}(x,y)$, computed using a 95% confidence level and a 2% margin of error with an average size of 2,337 reviews. This large sample is used for topic extraction to ensure the coverage of the user feedback space. Second, we draw a small sample denoted $S_{\text{small}}(x,y)$ from our review dataset, computed using a 95% confidence level and a 10% margin of error with an average size of 96 reviews. This sample is used for topic assignment, where manual evaluation is feasible and interpretable.
Figure 5. Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$. Insert category name is where we insert the name of the app category $y$. Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
Topic extraction. Because prior work reported mixed effects of few-shot examples (Wang_Yang_Wei_2023; Prakash_Wang_Hoang_Hee_Lee_2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023; Maram_LLM_Cure; Pham_Hoyle_Sun_Resnik_Iyyer_2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instructions with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
To mitigate hallucination during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention. To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
Figure 6. Template of the P5: topic assignment prompt used to assign topics to a list of reviews. We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Maram_LLM_Cure; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025).
To mitigate hallucination during topic assignment, we constrain the model’s output to a predefined set of deterministic labels, i.e., the extracted topics $T(x,y,n)$. The model must choose from these labels or select “Other” if none accurately represent the review. This closed-set format enables immediate detection of hallucination, as any label outside the predefined list signals a deviation from valid outputs. Additionally, structured CSV-formatted output further enforces adherence to the expected schema and prevents free-form or unconstrained generation. To ensure domain relevance, the five few-shot examples are handpicked from real reviews of Gen-AI apps, representing common contexts and linguistic patterns found in this domain (e.g., user reviews of AI Performance, Utility & Use Cases, or Content Quality).
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (cohen1960kappa) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labeling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan_Walmsley_1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann_Whitney_1947) with Bonferroni correction to adjust for multiple testing (3b6386c9-eeae-33a7-8c59-c0604aa87bef). We also use one-way ANOVA (Judd_McClelland_Ryan_2017) to test for significant differences across cleaning stages and few-shot settings.
Table 3. Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video). (Representative subset considered).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 86% 77% 91% 69% 95% 83% 84% 85% 85% 80% 90% 85% 72% 86% 77% 86% 89% 82%
E2: 3 shot 91% 86% 80% 66% 95% 83% 91% 80% 86% 82% 95% 87% 76% 83% 88% 92% 90% 86%
E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
4.2. RQ1 Findings
Observation 1.1: Providing more examples significantly improves classification and assignment accuracy. The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3. Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples. The Kruskal–Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$). Post hoc Mann–Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not. These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy. As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3). For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3. In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3. This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3. Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy. As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video). (Representative subset considered).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 83% 76% 90% 88% 76% 83% 84% 83% 89% 74% 82% 82% 81% 91% 81% 86% 89% 86%
E2: 3 shot 87% 91% 87% 86% 76% 85% 93% 78% 89% 89% 94% 88% 81% 88% 89% 91% 90% 88%
E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments. Given the inconsistencies observed in overall accuracy explained in observation 1.2, we isolate informative reviews to better evaluate model performance. Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3. This suggests that data cleaning improves the model’s ability to assign topics to reviews with informative content correctly. We conduct a one-way ANOVA test to evaluate whether the cleaning stage has a significant effect on classification accuracy. The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot. Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model. Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
Observation 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified. Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label). Most errors occurred in short or ambiguous reviews with limited context, where topic boundaries were less clear. These findings indicate that misclassifications are infrequent and stem from understandable edge cases rather than uncontrolled generation or hallucination, reinforcing the reliability of the LLM-based labeling process.
Summary of RQ 1 Our findings indicate that increasing the number of few-shot examples in the prompt significantly improves output accuracy. Additionally, filtering out non-informative reviews enhances the accuracy of extraction and assignment for informative reviews. The model using the 5-shot prompt combined with LLM-based filtering of non-informative reviews achieves the highest overall accuracy of 91%. Accordingly, we adopt this configuration for topic extraction in RQ2.
5. RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
Having validated the accuracy of our LLM-based topic extraction and assignment framework in RQ1, we now apply it to the full dataset to examine what users discuss in their reviews of Gen-AI apps. Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
5.1. Experiment Setup
In RQ1, we experiment with varying levels of data cleaning and numbers of few-shot examples. We find that filtering out non-informative reviews using an LLM and providing five-shot examples improves topic extraction and assignment accuracy. In RQ2, we apply the optimal topic extraction prompt identified in RQ1 to the remaining app categories and assign topics to all reviews in Stage 3. In this section, we provide a detailed description of the implementation of RQ2.
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totaling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. All reviews originate from Gen-AI apps; the distinction here is between review types, not between Gen-AI and non-Gen-AI apps. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Maram_LLM_Cure; Pagano_Maalej_2013; Chen_Lin_Hoi_Xiao_Zhang_2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform large-scale quantitative analysis on all automatically labeled reviews across both app categories and topic categories, followed by qualitative analysis on a representative sample. Specifically, for the quantitative component, we compute the following metrics:
  1. (1)
    Average Gen-AI Rating (AGR): the average star rating of reviews labeled with Gen-AI topics.
  2. (2)
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labeled with non-Gen-AI topics.
  3. (3)
    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI topics.
  4. (4)
    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics.
  5. (5)
    Gen-AI Review Percentage (GRP): the proportion of Gen-AI reviews relative to the total number of reviews.
  6. (6)
    Non-Gen-AI Review Percentage (NRP): the proportion of non-Gen-AI reviews relative to the total number of reviews.
We conduct the Mann–Whitney U test to examine whether reviews labeled as Gen-AI-related receive significantly higher ratings than those that are not. For developer-response analysis, we compute reply ratios, response delays, and response-type distributions for both Gen-AI and non–Gen-AI reviews, and we compare per-app reply behavior using a Wilcoxon signed-rank test (wilcoxon1992).
We complement our quantitative analyses with qualitative examination by manually coding statistically significant samples (95% confidence level, 10% margin of error) of both user reviews and developer responses. For user reviews we draw samples from every Gen-AI topic category resulting in a total of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
For developer responses, we perform two qualitative analyses. First, we code a representative sample of 96 replies drawn across all app categories to characterize the overall distribution of developer responses. Second, we manually code an additional 281 replies from three focal categories (Games, Education, and Entertainment) to capture deeper, category-specific response practices. Together, these qualitative analyses provide contextual grounding for the quantitative findings and offer a more complete understanding of both user concerns and developer engagement within Gen-AI apps.
5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories of Gen-AI apps. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality (i.e., AI Performance and Utility & Use Cases) exceeds the average rating of non-Gen-AI reviews (i.e., Technical Difficulties and Monetization Methods & Structure) in every app category of Gen-AI apps. This indicates that, when users explicitly comment on Gen-AI features, they tend to rate their experiences more positively than when they comment on other aspects of the same apps.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups (U=65,113,513,850U=65{,}113{,}513{,}850, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
This pattern is especially pronounced in categories such as Education and Books & Reference, which show some of the largest gaps between AGR and ANR. In Books & Reference, Gen-AI topics such as AI Performance and Human AI Interaction & Emotional Connection receive high ratings, while non-Gen-AI topics like Technical Difficulties and Customer Support are rated much lower due to crashes, generation failures, and unresolved support issues. Similarly, in Education, users highly rate Gen-AI topics such as Content Quality and Utility & Use Cases, but express dissatisfaction with Monetization Methods & Structure and Technical Difficulties. Together, these results indicate that strong Gen-AI features can substantially boost user satisfaction, but poor reliability, monetization, or support can still pull overall ratings down. Developers should therefore treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
Table 5. Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics (Full dataset considered).
App Category AGR ANR GRC GRP # Gen-AI Topics
Productivity 4.2 3.5 84,122 50% 6
Entertainment 4.0 3.2 56,672 48% 6
Education 4.6 3.3 31,465 75% 6
Art & Design 3.4 3.2 26,585 21% 3
Photography 3.9 3.3 15,567 12% 2
Tools 4.0 3.4 7,782 22% 3
Health & Fitness 4.8 4.0 6,896 88% 7
Video Players & Editors 4.6 3.9 5,179 14% 2
Games 3.6 2.7 1,788 35% 5
Music & Audio 3.0 2.6 1,679 18% 4
Books & Reference 4.1 2.3 428 39% 4
Average 4.0 3.2 21,651 38% 4
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As shown in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. Categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively, indicating that users in these domains directly engage with AI features to achieve personal or task-oriented goals such as self-improvement, learning, or workflow assistance.
Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category (Full dataset considered).
In contrast, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews, where 58% of feedback still focuses on non-Gen-AI topics such as usability, technical issues, and subscription or pricing models. This suggests that in creative or media-focused apps, users remain more concerned with traditional performance and value-for-money factors than with AI-driven functionality.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting how consistently users discuss Gen-AI features within each domain. Categories such as Entertainment and Health & Fitness exhibit the widest spreads, suggesting substantial variation among apps; some are highly AI-centric, while others receive minimal user attention to Gen-AI features. In contrast, categories like Video Players & Editors and Photography show narrow spreads concentrated at low percentages, indicating that across these domains, Gen-AI features consistently play an assistive or complementary role rather than being the central focus of user experience. Together, these patterns suggest that while certain app categories demonstrate uneven user emphasis on Gen-AI, others achieve relatively uniform user engagement of Gen-AI features.
Observation 2.3: Users discuss a wide range of Gen-AI topics, revealing where developers excel and where improvement is needed. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. Developers should pay particular attention to topics that appear most frequently across apps because they represent the aspects users discuss and evaluate most. Excelling in these areas is critical for maintaining competitiveness and user satisfaction. The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. Similarly, topics with high average ratings (e.g., Human AI Interaction & Emotional Connection and Utility & use cases) serve as benchmarks for best user experiences that other apps can emulate. Conversely, lower-rated or frequently criticized topics (e.g., Content Policy & Censorship and Content Quality) highlight recurring weaknesses across Gen-AI apps. Developers should treat these as priority areas for refinement to stay ahead of the competition and sustain user trust. We further examine each of the Gen-AI categories in the subsections below.
Table 6. Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
Gen-AI Topic Category GRC AGR # Apps # App Categories
AI Performance 61,071 4.0 127 10
Utility & Use Cases 56,505 4.8 45 3
Content Quality 44,594 3.6 131 8
Content Policy & Censorship 27,521 2.9 76 5
Human AI Interaction & Emotional Connection 25,142 4.8 13 4
Features & Functionality 16,275 4.0 58 4
Customer Support & Community 3,361 4.7 12 3
Comparison to Other Apps 2,315 3.9 13 3
Personalization & Customization 1,283 4.0 30 3
Accessibility & Inclusivity 96 4.1 4 1
Table 7. Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
Technical Difficulties 99,621 2.2 157 10
Monetization Methods & Structure 85,756 2.3 159 11
User Interface & Experience 80,827 4.5 156 11
Utility & Use Cases 21,464 4.8 34 1
Updates & Evolution 21,100 2.8 92 7
Personalization & Customization 6,860 4.8 18 1
Customer Support & Community 6,848 1.8 85 6
Limitations of Usage 5,971 2.5 45 4
Features & Functionality 1,810 3.0 11 1
Comparison to Other Apps 1,493 3.4 17 1
Accessibility & Inclusivity 401 2.6 5 2
Other (unassigned) 105,752 4.2 165 11
1) AI Performance. AI Performance refers to how well the AI model performs in terms of 1) accuracy (27%), 2) speed (22%), 3) understanding (15%), 4) memory (5%), 5) reliability (5%), and 6) creativity & randomness (3%). AI accuracy describes the precision of the AI response to the user’s prompt. AI speed represents how quickly the AI responds to a user’s prompt. AI understanding indicates the AI’s ability to comprehend the meaning, intent, and context behind the user’s prompt. AI memory signifies how well the AI remembers prior conversations and links them to the current one. AI can have both short-term memory, which spans a single conversation, and long-term memory, which spans multiple conversations. Reliability refers to the consistency and dependability of the AI system in performing its intended tasks without crashing, freezing, timing out, or failing to generate output. Creativity captures the degree of originality, uniqueness, or unpredictability in the AI’s responses.
Figure 8. Examples of reviews discussing AI Performance
Insights. We find that 79% of the users who discussed AI Performance provided a positive review with a rating of 4 or 5. Reviews 1 and 2 in Figure 8 show examples of positive reviews. Despite this overall satisfaction, 16% of reviews mentioned limitations in AI Performance, with the most frequently cited issue being AI understanding (5%), followed by inconsistency in accuracy (4%), reliability (3%), speed (2%), and AI memory (1%).
A recurring theme in the less favourable reviews is poor AI understanding, which leads to irrelevant or partially correct responses. Some users show awareness of prompt engineering and attempt to guide the AI with examples, yet still encounter issues. For instance, review 3 in Figure 8 illustrates a user struggling to communicate effectively with the AI, despite providing clear examples. Similarly, review 4 in Figure 8 highlights inconsistent performance for complex requests. While the AI excels at handling simple tasks, it often struggles with more complex ones.
Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
Figure 9. Examples of reviews discussing Content Quality
Insights. In our sample, 67% of reviews express satisfaction, praising the creative and polished outputs such as producing “radio-ready” songs and appending useful resources to textual answers as illustrated in reviews 1 and 2 in Figure 9 respectively.
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk_Siemenn_Surbehera_Chin_Tyser_Hunter_Raghavan_Hicke_Plummer_Kerret_et_al_2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial driver of overall user satisfaction and strategically incorporate these features to remain competitive. However, they should treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing. Even the most advanced AI functionality cannot compensate for poor technical reliability, weak customer support, or frustrating monetization structures. While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions. Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
3) Content Policy & Censorship. Content policy refers to rules and guidelines that define what type of AI-generated content is permissible. Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those rules.
Figure 10. Examples of reviews discussing Content Policy & Censorship
Insights. In our review sample, 73% of users comment on censorship experience. Of these 23% report frustration with the filtering that often block benign content, adversely affecting usability. For example, the author of review 1 in Figure 10 reports getting censored for saying “hi.” On the other hand, 10% of the raised concerns relate to lack of effective filtering such as unethical imagery involving children bypassing filters as expressed in review 2 in Figure 10. Regarding adult content, 67% of concerned users express dissatisfaction with the excessive restriction of access. While 19% explicitly request an optional censorship toggle for adult users as illustrated in review 3 in Figure 10. Additionally, users express concern over unclear content ownership. Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
$\Rightarrow$ Implication 2.
Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content. Ideally, developers could offer optional and adjustable censorship settings for mature or context-specific use. At the same time, platforms must clearly state who owns AI-generated outputs and how they can be reused, to maintain transparency and protect user trust.
4) Features & Functionality. Features & Functionality refer to the Gen-AI-driven capabilities offered by the app, such as voice interaction and prompt autocompletion, which shape the user experience and determine which features users find valuable or lacking.
Insights. In our sample, 75% of user reviews praise specific Gen-AI features, while 25% criticize them. For example, some users react negatively when a “stop-generation” button is removed and replaced with a voice mode feature, as illustrated in review 1 in Figure 11. While 11% of users enjoy the voice features, some describe them as useless. This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
Figure 11. Examples of reviews discussing Features & Functionality
5) Utility & Use Cases. Utility & Use Cases refer to the practical roles Gen-AI plays in users’ daily lives, including enhancing accessibility, supporting professional tasks, and enabling creative expression across diverse contexts.
Figure 12. Examples of reviews discussing Utility & Use Cases
Insights. In our sample, 64% of reviews emphasize AI’s educational utility, such as helping with homework, essay writing, tutoring, and breaking down complex topics as shown in review 1 in Figure 12. Additionally, some users specifically praise how Gen-AI can act as an accessible on-demand learning assistant for dyslexic users, as commented in review 2 in Figure 12. Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
6) Personalization & Customization. Personalization & Customization describe how Gen-AI systems adapt to user preferences and allow users to tailor outputs. This includes customizing output formats, editing generated content, and adopting preferred AI personas, e.g., voices or AI character roles.
Figure 13. Examples of reviews discussing Personalization & Customization
Insights. In our sample, 38% of reviews emphasize the AI’s inability to create a personalized experience. For example, review 1 in Figure 13, discusses the app’s inability to maintain shopping and to-do lists. Additionally, 10% of users request inline editing capabilities to support customization by refining generated content directly, rather than having to regenerate entirely. For instance, review 2 in Figure 13 suggests the option to edit lyrics in an already generated song rather than regenerate a new one entirely. Meanwhile, 9% of reviews request more diversity in AI characters and custom voices. This reflects a growing interest in personalizing AI identities to match preferences such as female voice options. For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
$\Rightarrow$ Implication 3.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences. Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized. Despite the launch of tools like ChatGPT Edu (openaiChatGPTedu), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
Figure 14. Examples of reviews discussing Comparisons to Other Apps
Insights. We find that 66% of the reviews draw comparisons to other AI-powered apps (e.g., comparing Microsoft Bing Search with Google). We also observe that 9% of the reviews compare the app with non-Gen-AI alternatives. For instance, review 1 in Figure 14 criticizes the forced integration of the AI chat feature, expressing a strong preference for traditional search functionality. Finally, 18% of reviews draw comparisons between the apps they review and general-purpose AI tools (e.g., ChatGPT), highlighting expectations shaped by user experiences with AI tools, as shown in review 2 in Figure 14. 77% of the reviews that compare the app with ChatGPT believe that ChatGPT outperforms the app in question, while 15% prefer the performance of the domain-specific app. These comparisons highlight a key challenge for domain-specific Gen-AI app developers. Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
8) Customer Support & Community. Customer Support & Community refers to reviews that address developers directly to express their experience, such as reporting user enjoyment with specific features or requesting to fix app shortcomings.
Figure 15. Examples of reviews discussing Customer Support & Community
Insights. In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
$\Rightarrow$ Implication 4.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary). User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Maram_LLM_Cure; PeerApps_SANER; Assi_Hassan_Tian_Zou_2021). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
9) Human AI Interaction & Emotional Connection. Human AI Interaction & Emotional Connection reviews highlight the emotional and relational dimensions of human–AI interaction, particularly with conversational Gen-AI apps.
Figure 16. Examples of reviews discussing Human AI Interaction & Emotional Connection
Insights. We observe that users enjoy their interactions with conversational Gen-AI apps for multiple reasons. First, users enjoy a realistic interaction (i.e., closely similar to human dialogue). This perceived realism fosters a sense of companionship and emotional connection (e.g., review 1 in Figure 16). Second, reviews highlight the comfort in the emotional support provided by Gen-AI apps, particularly because the interaction lacks real human judgment, as shown in review 2 in Figure 16. This illustrates the importance of AI’s perceived neutrality and non-judgmental nature, especially when discussing sensitive or emotionally charged topics. Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
10) Accessibility & Inclusivity. Accessibility and inclusivity refer to the degree to which the Gen-AI apps accommodate users for affordability, multilingual support, age appropriateness, and usability for individuals with cognitive or physical impairments.
Figure 17. Examples of reviews discussing Accessibility & Inclusivity
Insights. Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
$\Rightarrow$ Implication 5.
The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability. Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care. Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
The findings further highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
Observation 2.4: Developer Responses to Gen-AI Reviews Are Primarily Courteous or Diagnostic, With Few Immediate Resolutions. Table 8 shows that for Gen-AI apps, the most common developer response type is Acknowledgment and Appreciation (31%), which consists largely of brief thank-you messages. These replies most often accompany positive reviews and serve as courteous engagement rather than substantive technical follow-up.
The second most frequent category, Request for Additional Information (30%), reflects the practical limits of troubleshooting Gen-AI issues through review threads. Even when space is not a constraint, users rarely provide the detailed context needed to diagnose problems that relate to model behavior, device configuration, or account-specific issues. Developers therefore redirect users to email or in-app support channels where logs, device information, or credentials can be shared securely.
By contrast, Issue Resolved and Update Advised appears rarely (2%), whereas Acknowledged and Under Investigation is more common (10%). This pattern aligns with the nature of Gen-AI systems: many issues require model adjustments, infrastructure changes, or updates to safety and moderation mechanisms, none of which can be addressed immediately or within the constraints of a review thread. As a result, developers are more likely to acknowledge problems and communicate that work is underway than to confirm the issue has been resolved.
Table 8. Summary of developer response types, including descriptions, number of reviews, and percentage distribution (Representative subset considered).
Topic Description # Reviews %
Acknowledgment and Appreciation Generic responses that thank users for their feedback, often including a brief note about considering future updates or improvements. 30 31%
Request for Additional Information Responses indicating that the reported issue lacks sufficient detail and requesting that the user contact the developer directly for further clarification or resolution. 29 30%
Clarification or Policy Rebuttal Responses that clarify or refute user concerns, often explaining app policies or correcting misunderstandings. Most common for reviews concerning free versus premium features. 13 14%
Guidance and Troubleshooting Tips Responses that address reported issues by offering troubleshooting advice, resources, or specific steps to help users resolve the problem. 12 13%
Acknowledged and Under Investigation Responses confirming that the development team is aware of the issue and is currently investigating or working on a solution. 10 10%
Issue Resolved and Update Advised Responses informing the user that the reported issue has been resolved and recommending that they update or reinstall the app. 2 2%
Total 96 100%
Hassan et al. (developer_response_safwat) analyzed developer replies across a broad set of mobile apps and identified recurring response drivers such as “thank the user”, “ask for more details”, “provide guidance”, “ask for endorsement”, “advertise other products”, wait for updates, and “justify advertisements”. Many of these practices align closely with our taxonomy. For example, their “thank the user” category maps to our “Acknowledgment and Appreciation,” their “ask for more details” category maps to our “Request for Additional Information,” and their “justify the advertisements” category maps onto our broader Clarification or Policy Rebuttal, which includes explanations about subscription models, premium features, and content restrictions. Notably, their work does not treat “issue resolved” as a dominant response driver, and we similarly observe that confirmed resolutions are relatively rare in our Gen-AI dataset (2%). We report the category explicitly for completeness of Gen-AI developer behaviors, but its low prevalence mirrors prior findings rather than contrasting with them. Overall, the strong correspondence between the two taxonomies suggests that general response practices extend to Gen-AI apps, while our domain-specific categories capture the additional emphasis developers place on clarifying AI-related behavior and managing expectations around the evolving nature of Gen-AI features.
Observation 2.5: App categories show highly distinctive developer response patterns. Table 9 reveals variation in how developers across different app categories respond to user feedback. Some categories exhibit exceptionally high engagement, while others show very limited interaction. For example, Games stands out with both the highest reply ratio (56%) and the longest average response delay (44.7 days). In contrast, Education demonstrates the shortest response times, with developers replying in less than a day on average (0.7 days). Meanwhile, Entertainment apps show the lowest reply ratio among categories with meaningful activity (2%). These contrasting patterns suggest that developer responsiveness is shaped not only by category size but also by the operational demands, user expectations, and support norms unique to each domain.
Table 9. Developer response activity across app categories, including total reviews, total developer replies, overall reply ratio, and average response delay (Full dataset considered).
Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
Productivity 169,119 8,407 5% 8.4
Photography 126,277 31,070 25% 4.5
Art & Design 125,362 39,593 32% 5.5
Entertainment 118,669 2,783 2% 1.7
Education 41,905 16,034 38% 0.7
Video Players & Editors 36,507 11,123 30% 2.2
Tools 35,046 905 3% 32.45
Music & Audio 9,150 2,127 23% 3.4
Health & Fitness 7,849 338 4% 7.7
Games 5,070 2,848 56% 44.7
Books & Reference 1,112 1 0% 1.0
All Categories 676,066 115,164 20% 10.2
Our manual analysis of developer responses in the Games category helps explain why it exhibits both the highest reply ratio and the longest response delay. Responses in which developers report that an issue has been fixed account for 30% of all replies in this category, far exceeding the 2% observed in the overall dataset. This emphasis on confirmed fixes aligns with the long response delay shown in Table 9. This likely reflects that developers often wait until a patch is deployed before responding, resulting in slower but more substantive communication. Responses indicating that an issue is acknowledged and under investigation represent 13% of replies, slightly higher than the overall 10%. Together, these categories show that Games developers place strong emphasis on technical issue resolution and ongoing investigation.
Generic thank-you messages appear in only 12% of replies, considerably lower than the overall 31%. The remaining responses highlight active engagement with user issues: 28% request additional information, 10% provide troubleshooting guidance, and 8% offer clarifications or policy rebuttals. Overall, developer communication in Games is markedly problem-oriented and resolution-driven, which leads to both high engagement and extended response timelines.
A closer look at Education apps helps explain why this category exhibits the fastest average developer response time. In our manual sample, the majority of Education responses fall under Acknowledgment and Appreciation (73%), which is markedly higher than both the overall rate of 31% and the 12% observed in Games. These replies are typically short, template-based thank-you messages that require minimal effort to generate, allowing developers to respond quickly and consistently.
Education apps also show a comparatively high rate of Request for Additional Information (16%). This proportion is notably larger than the overall sample and reflects a common pattern in Gen-AI Education reviews. Users frequently report that the app produced incorrect answers, yet they often omit the original prompt, the model’s response, or screenshots of the interaction. Without these details, developers cannot diagnose or reproduce the issue within the review thread. As a result, they quickly reply by directing users to contact support channels where complete context can be shared securely. This response type is therefore both common and fast, since it does not require problem investigation within the review platform itself.
By contrast, response types requiring deeper engagement occur far less frequently in Education, including Guidance and Troubleshooting Tips (2%), Acknowledged and Under Investigation (6%), and Issue Resolved and Update Advised (1%). Together, these patterns show that Education apps tend to rely primarily on quick acknowledgments or requests for additional information, which contributes to the category’s exceptionally short response delay.
Entertainment apps show the lowest reply ratio in our dataset, and the distribution of response types helps explain this pattern. Acknowledgment and Appreciation responses account for only 14%, far below the overall rate of 31%. This indicates that developers in this category do not typically reply to routine or positive reviews and instead reserve responses for situations where user action or developer follow-up is required. Correspondingly, a large share of replies fall under Request for Additional Information (38%) and Guidance and Troubleshooting Tips (22%), suggesting that developers primarily intervene when users report crashes, playback issues, or account problems that require diagnostic detail. The relatively high proportion of Acknowledged and Under Investigation (19%) further shows that many of these reports involve technical issues that cannot be resolved immediately, which may discourage broader engagement with lower-priority feedback. Together, these patterns indicate that Entertainment developers focus their limited response efforts on actionable problem reports rather than general user engagement, resulting in a lower overall reply ratio.
Observation 2.6: Developer responsiveness varies substantially across Gen-AI topic categories. Table 10 summarizes developer reply behavior for each Gen-AI topic category. The topic with the highest reply ratio is Accessibility & Inclusivity (47%), despite having the smallest number of reviews. This indicates that developers treat accessibility-related feedback as a high-impact area that warrants immediate and consistent attention. Such reviews often report barriers to use for specific user populations, making them critical from both ethical and product-quality perspectives. The short average delay (0.2 days) further reflects the urgency developers assign to this category.
Beyond accessibility, the next highest reply ratios are found in Utility & Use Cases (18%) and Content Quality (14%), suggesting that developers are highly responsive to feedback about how users apply AI features and whether the generated outputs meet expectations. These topics directly affect user satisfaction and daily usability, which likely motivates more frequent developer engagement.
In contrast, the categories with the lowest reply ratios are Customer Support & Community (1%), Human–AI Interaction & Emotional Connection (2%), and Content Policy & Censorship (5%). These topics often involve complaints about communication quality, emotional tone, safety constraints, or moderation decisions. Such issues are typically systematic, tied to model behavior or policy rules that cannot be quickly adjusted, which may explain the lower response frequency and the longer average delays seen in these categories.
This contrast between highly actionable categories and structurally constrained categories reinforces the idea that responsiveness is shaped by whether an issue is immediately fixable or requires deeper system changes. Overall, these patterns show that developers prioritize issues that are actionable in the short term (e.g., accessibility barriers, inaccurate outputs, task-level failures) while responding less frequently to concerns that require deeper model changes or policy revisions. This reflects a practical triaging strategy driven by the nature of Gen-AI systems and the feasibility of immediate fixes.
Table 10. Developer response patterns across Gen-AI topic categories, including total reviews, total developer replies, reply ratio, and average response delay (Full dataset considered).
Topic Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
AI Performance 61,071 7,152 12% 4.2
Utility & Use Cases 56,505 9,895 18% 0.6
Content Quality 44,594 6,275 14% 4.3
Content Policy & Censorship 27,521 1,397 5% 7.9
Human–AI Interaction & Emotional Connection 25,142 542 2% 7.5
Features & Functionality 16,275 1,892 12% 5.5
Customer Support & Community 3,361 28 1% 10.3
Comparison to Other Apps 2,315 287 12% 6.0
Personalization & Customization 1,283 106 8% 2.1
Accessibility & Inclusivity 96 45 47% 0.2
All Topics 238,168 27,619 12% 4.9
Observation 2.7: Developers respond less frequently to Gen-AI reviews than to non-Gen-AI reviews as illustrated in Figure 18. We compared per-app developer response rates for Gen-AI and non-Gen-AI reviews using a paired non-parametric Wilcoxon signed-rank test ($n=164$ apps). Developers replied to a median of 9.6% of Gen-AI reviews versus 17.7% of non-Gen-AI reviews. Although the median per-app difference is modest (0.7-0.7 percentage points), the effect is statistically significant ($W=2911$, $p<0.001$), indicating a consistent pattern across apps.
This gap aligns with the qualitative patterns observed earlier. Gen-AI reviews often describe failures rooted in model behavior, data quality, or moderation mechanisms, which are difficult to diagnose from short review text. By contrast, non-Gen-AI reviews commonly involve UI issues, feature requests, or account problems that are easier to acknowledge, triage, or resolve within the constraints of the review interface. As a result, developers appear slightly more willing to engage with non-AI feedback where their replies can directly influence user experience. In addition to analyzing user perspectives, we also examine how developers respond to Gen-AI reviews, capturing patterns of engagement, resolution, and support practices across app categories.
Observation 2.8: Developers respond to Gen-AI and non-Gen-AI reviews at similar speeds. Figure 18 compares developer response delays for Gen-AI and non-Gen-AI reviews across all apps. While Gen-AI reviews show a slightly lower median delay (a difference of 0.15-0.15 days), this difference is not statistically significant. In practical terms, once developers choose to reply, they tend to respond at comparable speeds regardless of whether the review discusses Gen-AI functionality or non-AI aspects of the app. This suggests that the lower engagement with Gen-AI reviews observed in Observation 2.7 reflects differences in response likelihood rather than response promptness.
Figure 18. Comparison of developer responsiveness across apps. Left: Relative likelihood of developer response to Gen-AI reviews compared to non–Gen-AI reviews, expressed as the ratio of per-app response probabilities (values below 1 indicate lower responsiveness to Gen-AI reviews). Right: Distribution of mean developer response delays per app for Gen-AI and non–Gen-AI reviews; the two delay distributions do not differ significantly. (Full dataset considered.)
$\Rightarrow$ Implication 6.
Taken together, our developer response findings show that review threads in Gen-AI apps primarily serve as an initial point of contact rather than a channel for resolving AI-specific failures. Developers face structural barriers when responding to Gen-AI feedback, since many issues require information that cannot be shared in public reviews or involve model- or policy-level changes that cannot be deployed quickly. As a result, developers tend to reserve detailed engagement for issues that are actionable in the short term and redirect users elsewhere when deeper diagnostic context is required.
These patterns highlight a broader platform-level limitation: while app stores surface user feedback effectively, they do not provide the depth of interaction required to investigate or clarify Gen-AI behaviours. There is therefore an opportunity for platform owners to introduce more structured, built-in support mechanisms that complement review sections and enable richer, privacy-preserving communication channels. Such improvements would better support both developers and users as Gen-AI apps become more complex and more central to mobile ecosystems.
Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents. They expect fast, creative, and context-aware responses, along with personalization and transparency. For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT). Our developer-response analysis reinforces this point: developers most often engage with actionable, user-reported issues and struggle to address Gen-AI failures directly within the review platform. This suggests that better support channels will be essential for meeting rising user expectations.
6. RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
To understand how users perceive Gen-AI features over time, we explore the temporal dynamics of user feedback and ratings in app reviews. Our goal is to uncover how user opinions about Gen-AI topics shift as Gen-AI capabilities mature and become more integrated into mobile apps.
6.1. Experiment Setup
We track the evolution of user feedback over time by conducting a two-level trend analysis. First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time. This perspective helps highlight the increasing user attention to Gen-AI features. Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns. We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis. Based on the observed rating dynamics, we select three representative topic categories for deeper investigation: Content Quality (decreasing trend), AI Performance (stable trend), and Content Policy & Censorship (increasing trend). The following steps outline how we conduct the analysis:
Step 1: data preparation. We group reviews according to their topic category $(T)$ and time period $(P)$. Specifically, for every topic category $T$, we organize reviews into half-year intervals (i.e., every 6 months) based on their submission dates. Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of reviews in period $P$ that belong to category $T$. We exclude the topic category Accessibility & Inclusivity because it only contains 96 reviews in total, which are not enough to conduct a time series analysis.
Step 2: clustering methodology. To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends. For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we extract slope and standard deviation as features following the three-step process listed below:
  1. 1.
    Features extraction. Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash_Koot_2009). Feature selection enhances interpretability, aligns with research goals  (Kreienkamp_Agostini_Monden_Epstude_DeJonge_Bringmann_2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable). For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$. Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend. It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
  2. 2.
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren_Liao_2005). We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023; 10.1145/3721125). K-Means is suitable for low-dimensional feature spaces (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
  3. 3.
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari_DwiNugroho_Baraku_NovriYanda_Liwardana_2023), across multiple values of $k$. Based on the silhouette results, we select the $k$ value closest to 1. The optimal silhouette scores correspond to kk=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$. We complement this with qualitative validation through manual inspection, confirming that kk=3 yields the most interpretable and distinct groupings.
Step 3: manual trend investigation. To interpret the identified clusters, we perform qualitative analysis on statistically significant random review samples from key time periods. For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline. From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error. We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations. This process provides qualitative context that enhances our understanding of the observed quantitative trends.
6.2. Findings
Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time. Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics. As illustrated in Figure 19, the percentage of reviews discussing Gen-AI topics increases over time. The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features.
This growth indicates that users are increasingly evaluating apps through the lens of their Gen-AI features. As AI becomes a central part of the user experience, developers can no longer treat it as an add-on; users expect it to perform reliably and meaningfully enhance the app’s core value. The rising share of Gen-AI reviews therefore signals that user satisfaction and engagement increasingly depend on how well these features are designed, integrated, and maintained.
Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 20 with only a slight change of ± 0.2 stars across the observed period. Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories. As illustrated in Figure 21, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable. These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings. The following observations examine these diverging topic-specific trajectories in greater detail.
Figure 19. Evolution of the Percentage of Gen-AI Reviews (Full dataset considered).
Figure 20. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics (Full dataset considered).
Gen-AI topics analysis. As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters. These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 21 and 22 for the average ratings and the percentage of reviews, respectively.
To deepen our understanding of the observed rating trends, we perform a manual trend investigation, i.e., a qualitative analysis, on three representative topic categories, each corresponding to a distinct trend direction. For the Decreasing trend, we select Content Quality due to the unexpected nature of its decline, particularly in light of ongoing advancements in AI capabilities as shown in Figure 21. For the Stable trend, we examine AI Performance, where we had anticipated an upward trend reflecting improvements in system performance. For the Increasing trend, we analyze Content Policy & Censorship to investigate how user perceptions of policy changes have evolved and to uncover the factors contributing to rising satisfaction in this area.
Figure 21. Trend Clusters for Average Ratings of Gen-AI Topics (Full dataset considered).
Figure 22. Trend Clusters for Percentage of Gen-AI Topics (Full dataset considered).
Observation 3.2: The Content Quality topic category shows a decline in rating over time, despite improvements in Gen-AI technology. As shown in Figure 21, the average rating for Content Quality decreases from 4.5 in 2022-H1 to 3.3 in 2024-H2. To better understand the unexpected downward trend in Content Quality ratings, we conduct a manual analysis of user reviews across three strategically selected time periods: 2022-H1 (corresponding to the highest average rating), 2022-H2 (marking a period of sharp decline), and 2024-H2 (representing the lowest observed average rating). This sampling allows us to investigate user behavior before, during, and after the major shifts in rating averages. Based on the manual inspection, we identify two key user-related factors driving the decrease as follows.
Figure 23. Examples of reviews discussing Content Quality at different time periods.
Observation 3.3: Over time, the Content Quality reviews become less tolerant of flaws and more likely to penalize them. Excitement for emerging Gen-AI technologies appears to influence early adopters’ evaluation criteria, making them less likely to penalize imperfections. In our analysis of the 2022-H1 sample, 96% of the reviews are positive, with only 5% of them including constructive feedback despite assigning a perfect 5-star rating. For example, the user writing review 1 as shown in Figure 23 awards the app the highest rating while reporting that the AI generated low-quality facial images. Such reviews reflect a pattern in which users prioritize innovation and novelty over technical quality, demonstrating a higher tolerance for flaws during the early adoption phase.
Over time, users become less tolerant of imperfections. The percentage of positive reviews declines from 96% in 2022-H1 to 59% in 2024-H2, indicating a significant shift in user ratings. Moreover, positive reviews become increasingly critical: in 2022-H2, 12% of all reviews offer constructive feedback while still assigning a high rating (4 or 5 stars), rising to 18% in 2024-H2. This trend suggests that users’ expectations have heightened.
More notably, an increasing number of users begin deducting rating star points for flaws over time. In 2022-H2, only 2% of constructive reviews assign a 5-star rating, while 10% assign a 4-star rating. By 2024-H2, this trend intensifies: 15% of constructive reviewers assign a 4-star rating, whereas only 4% assign a 5-star rating. For instance, review 2 in Figure 23 expresses the need for improvement in graphics and gives a 4-star rating. This shift presents a growing criticality among users. While early reviews tend to overlook shortcomings in favour of innovation, later reviews are more discerning, penalizing points even when expressing generally positive sentiment.
Observation 3.4: Over time, the Content Quality reviews are more experimental and demanding. Initial reviews frequently convey surprise and satisfaction with relatively modest capabilities, indicating limited expectations during the early adoption phase. For example, review 3 in Figure 23 from 2022-H1 expresses surprise at the high quality of the content, despite initially having low expectations. Such statements suggest that early positive ratings are often driven by low expectations of the novel technology rather than its actual performance.
As Gen-AI technologies become more widespread, user expectations rise accordingly. Reviews begin to reflect more ambitious and exploratory usage patterns, with users increasingly testing the system’s capabilities through complex prompts. When the app fails to meet these heightened demands, users express dissatisfaction. For example, review 4 in Figure 23 in 2022-H2, reports success in simple requests but horrific results with complex requests, ultimately giving the app a poor 1-star rating. This evolution in review content presents a shift from initial awe to critical evaluation, as users become more experimental and less tolerant of limitations.
We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar_2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity. Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino2024). Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1. To account for this overall stability, we analyze fluctuations at the app-category level. We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others, producing the flat global average.
However, this apparent stability conceals divergent user experiences. Categories such as Education and Games show clear improvement, reflecting that users in these domains perceive AI as directly enhancing learning and task performance. In contrast, most other categories (e.g., Productivity, Entertainment, Books & Reference) show gradual declines. This may suggest that developers in these areas are integrating AI features prematurely, before the technology is fully optimized for their use cases. When users accustomed to fast and accurate AI systems encounter slower or less reliable integrations, their satisfaction drops. These declines therefore point to gaps in seamless integration and contextual adaptation rather than outright failure of AI Performance.
For developers, this indicates that AI performance has become a key competitive factor. Users now expect fast, accurate, and context-aware AI behavior. When AI integrations fall short or introduce lag, inconsistency, or irrelevant responses, users quickly express dissatisfaction. Sustaining trust therefore depends on ensuring that AI features operate seamlessly within the app’s core experience rather than as add-ons.
Figure 24. Examples of reviews discussing AI Performance at different time periods.
Observation 3.6: Advances in image generation capabilities contribute to the slight increase in the average rating of the AI Performance topic in 2023-H1. In 2022-H1, negative reviews of image generation apps constitute 9% of reviews in our sample and account for 47% of all negative reviews during that period. These users express frustration with the visual output. For example, review 1 in Figure 24 expresses frustration over the performance of the image generator.
By 2023-H1, these issues appear to be addressed, as the percentage of negative reviews of image generation apps drops to 5%, and their share of total negative reviews decreases to 29%. This reduction in dissatisfaction likely suggests improvements in the AI Performance of image generation, which helps lift the overall AI Performance rating during this period.
Observation 3.7: Underperformance in Productivity apps leads to a slight decline in the average rating of the AI Performance topic in 2024. By 2024-H1, Productivity apps emerge as the leading source of performance-related complaints, accounting for 8% of all reviews in our sample and 44% of all negative reviews. Unlike creative tools, which are often evaluated in terms of novelty or expressiveness, Productivity apps are frequently judged against well-established, non-AI alternatives that offer straightforward functionality. Consequently, users express notable frustration with AI-driven features that are perceived as unnecessarily complex or unreliable. For example, review 2 in Figure 24 prefers a regular assistant over an AI assistant because the AI assistant complicates simple tasks, and review 3 in Figure 24 prefers a calculator over an AI due to accuracy issues.
These comments highlight that Gen-AI integration can complicate tasks that would otherwise be simple. Rather than enhancing the user experience, AI features may undermine usability if they are poorly aligned with user expectations or fail to deliver practical value. Some reviews explicitly question the necessity of AI features. Review 2 in Figure 24 describes abandoning their AI assistant after it offered a detailed classification of fan technologies instead of simply turning on a fan. This indicates that forced integration may backfire, especially when it disrupts core functionality.
Observation 3.8: The stability in the average rating of the AI Performance topic is driven by offsetting trends across app types. Some app categories show improvement (e.g., image generation), while others show decline (e.g., Productivity). These shifts tend to cancel each other out at the global level. In 2023, rising satisfaction with creative apps contributes to a mild increase in performance ratings. In 2024, dissatisfaction in Productivity apps is balanced by relatively stable perceptions in creative and educational apps. This compensatory effect explains the flat trend in average performance ratings over time.
Although the global average for AI Performance ratings remains stable over time, this apparent stability masks meaningful variation across app categories. Our analysis reveals divergent trends: categories such as Art & Design, Entertainment, and Books & Reference show moderate declines in average AI Performance ratings (–0.84 to –0.93), while others like Education and Games experience notable increases (+1.39 and +0.96, respectively). These opposing shifts suggest that stability at the aggregate level may result from the balancing of improvements and declines across distinct domains. This trend may also reflect differences in user expectations. Creative tools are often evaluated for imagination and output quality, while Productivity apps are judged more on efficiency and reliability.
$\Rightarrow$ Implication 7.
Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users. Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust. AI integration should be purpose-driven and context-aware. Raji et al. (Raji_Kumar_Horowitz_Selbst_2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility. The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration. Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default. Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1.
To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1. We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
This pattern reflects how content policy concerns were especially prominent in the early stages of Gen-AI deployment, when moderation systems were still limited and users frequently encountered unsafe or poorly filtered outputs. Over time, developers appear to have become increasingly aware of these issues and implemented more effective safeguards, leading to gradual improvement in user satisfaction. The upward trajectory in ratings thus signals a maturing ecosystem where content safety and moderation mechanisms are continuously refined. For developers, this emphasizes the importance of iterative moderation design. Continuous refinement of content filters not only improves safety but also builds long-term user trust.
Figure 25. Examples of reviews discussing Content Policy at different time periods.
Observation 3.10: Ethical concerns around AI-generated art are more prominent in early samples but disappear in later ones. In 2022-H2, 10% of sampled reviews express concerns about the ethics of training AI on copyrighted human art without consent or compensation. For example, review 1 in Figure 25 expresses how Gen-AI harms artists whose art is stolen. This concern disappears from the random samples of 2023-H2 and 2024-H1.
This highlights the importance of transparent and ethical data use policies. However, the absence of similar comments in later samples suggests either that the user-base shifted or that these concerns were deprioritized in users’ feedback. The drop may also indicate user desensitization or limited platform changes aimed at addressing the issue. Nevertheless, the ethics of AI content policy should still be addressed by researchers and developers.
Observation 3.11: Content filters improve over time in blocking unwanted adult content (i.e., the raised concerns drop from 12% to 3%) and illegal content (i.e., the raised concerns drop from 4% to 0%). In our early sample, 15% of the studied reviews flag harmful content slipping past moderation, especially involving sexual or predatory material. This drops to 5% in 2023-H2 and to 3% by 2024-H1, indicating increased user satisfaction with content safety mechanisms. Examples of such reviews can be found in reviews 2, 3, and 4 in Figure 25.
The proportion of users reporting the unintentional generation of adult content drops from 12% in 2022-H2 to 3% in 2024-H1. Reports of illegal or predatory content involving children drops from 4% to 0%. These declines reflect measurable improvement in censorship mechanisms for harmful content.
Observation 3.12: Complaints about false-positive censorship grow over time (from 35% to 59%), especially among users engaging in regular and appropriate use. As filters improve at blocking harmful content with time, users increasingly report unnecessary censorship during benign interactions. In 2022-H2, 35% of users in our sample request a less strict filter. This jumps to 71% in 2023-H2 and settles at 59% in 2024-H1.
While the proportion of users requesting more freedom to generate explicit or violent content decreases slightly (from 21% in 2022-H2 to 14% in 2024-H1), more users begin to complain about filters interfering with regular usage. For example, 6% of reviews in 2022-H2 flag such issues, rising to 30% in 2024-H1. Review 5 in Figure 25 shows an example of reviews discussing false-positive censorship, where the user reports receiving a guidelines warning for writing a non-explicit role-play scene. This suggests that while filters are more successful in removing harmful content, they are over-correcting and obstructing benign, creative, or functional uses. Over-correction can hinder the Gen-AI user experience, particularly in role-play and storytelling apps.
Observation 3.13: Despite an overall positive trend, over-correction in censorship filters causes a minor ratings setback in 2024. The average rating for Content Policy and Censorship follows an increasing trend over time, reflecting growing user trust in content moderation systems. As previously discussed, this trend is supported by a substantial decline in complaints about harmful content. Correspondingly, the share of positive reviews that discuss censorship rises from 4% in 2022-H2 to 51% in 2023-H2. This rise indicates that while many users enjoy the app enough to leave a high rating, they still express frustration with the limitations imposed by the censorship filters.
However, this progress introduces a new challenge of false-positive filtering in 2024-H1. As previously discussed, while moderation systems are successful in blocking inappropriate content, they begin flagging benign prompts as violations. This over-correction results in a moderate decline in the share of positive censorship-related reviews (20% in 2024-H1) and contributes to a slight drop in the average rating. Nevertheless, this dip does not reverse the overall gains. Ratings in 2024-H1 remain higher than in 2022-H2, suggesting that while usability concerns emerge, users still view the improvements in content safety as a net positive. This also highlights an important reality for developers; progress in content moderation is unlikely to be perfectly linear. Minor setbacks and fluctuations are to be expected as systems evolve and adapt to user needs.
$\Rightarrow$ Implication 8.
Developers must strike a careful balance between safety and creative freedom. While stricter filters reduce the occurrence of disturbing or illegal content, they increasingly disrupt legitimate use. To avoid over-correction, developers should implement modular moderation systems with customizable filters or toggles that can be adjusted by personal preference. Clear content policy guidelines and transparent communication about AI behaviour are also essential. Additionally, ethical considerations around AI training data should not be ignored by developers or researchers, even if fewer users express concerns. These considerations remain essential for maintaining long-term trust and ensuring platform responsibility.
Summary of RQ 3 Our analysis reveals that Gen-AI topics evolve differently over time with increasing, stable, or decreasing trends. Content Quality ratings steadily decline, not due to technological regression, but due to increased user expectations. In contrast, AI Performance shows a stable rating trend, driven by a balancing effect between improving and declining domains. Lastly, Content Policy & Censorship follows an overall positive trend, with improved content filters increasing user ratings. These findings reflect a maturing user base that is more experienced, discerning, and vocal. Our user base of Gen-AI apps raise the bar for what constitutes a high-quality Gen-AI experience.
7. Discussion
In this section, we reflect on the broader implications of our findings and provide targeted recommendations for six key stakeholder groups: 1) users, 2) platform owners, 3) developers, 4) testing tool designers, 5) requirements engineers, and 6) policymakers and regulators.
7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 1, 3, and 5). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 5).
7.2. Implications for Platform Owners
Platform owners must bear responsibility for ensuring that Gen-AI apps comply with ethical standards (Observation 3.10). As AI capabilities evolve rapidly, app store operators should implement both technical tools and regulatory frameworks that evaluate whether submitted Gen-AI apps meet baseline criteria for safety, transparency, and responsible data usage. This may include requiring apps to disclose whether AI is used, clarifying how outputs are generated, and verifying that training data sources are ethically obtained. By introducing certification processes or audit mechanisms for Gen-AI features (particularly in apps that claim educational, therapeutic, or professional utility), platforms can promote accountability and mitigate harm, fostering greater trust among users and developers alike.
App platforms play a critical role in shaping the quality of user feedback. Our results show that a large proportion of user reviews are non-informative (Observation 1.2), which limits their value for developers and researchers. Future research could explore how redesigning review interfaces might encourage more meaningful feedback. One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
Beyond improving review quality, our findings also reveal structural limitations in how app stores currently support developer–user interaction. As shown in our developer response analysis (Implication 6), review threads in Gen-AI apps function primarily as an initial point of contact rather than a space for resolving AI-specific failures. Many issues require diagnostic information that cannot be shared publicly or involve model- and policy-level changes that cannot be deployed quickly. Platform owners can play an important role by offering richer, privacy-preserving support channels that complement review sections and allow developers to investigate Gen-AI issues more effectively. Strengthening these communication pathways would improve user experience and reduce the friction developers face when addressing complex AI behaviours.
7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 4). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 3 and 5). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 2), supporting accessibility and inclusivity concerns (Implication 5), and enabling more personalized and flexible user experiences (Implication 3). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 7).
7.4. Implications for Testing Tool Designers
Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Observation 1.4), demonstrating the effectiveness of prompt design and data pre-filtering (Observations 1.1 and 1.3). These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows. Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment. Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles.
7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis framework provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality. Beyond Gen-AI apps, the same framework and analytical tools can be applied to user reviews in other domains, enabling engineers to systematically identify evolving requirements and usability concerns across a wide range of software systems.
7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 2). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 2), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
8. Threats to Validity
Construct threats to validity refer to the risk that the measures we use in our study do not accurately reflect the concepts we aim to study. First, we use the average app review rating as a proxy for user satisfaction. However, this metric may not fully capture the sentiment expressed in the review text. For example, users sometimes assign high ratings while simultaneously reporting critical issues or limitations. To mitigate this threat, we supplement the quantitative rating data with manual qualitative analysis of review content, ensuring that nuanced sentiments, both positive and negative, are appropriately captured. Second, our methodology relies on LLMs to automatically filter and classify reviews. To control for potential biases introduced by the LLM-based automation, we manually annotate samples and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks. Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data. A key internal threat relates to the uncertainty surrounding the timeline for introducing the Gen-AI features app. While we manually verify that all selected apps currently support Gen-AI functionality, we cannot determine the exact date of integration. To reduce the risk of analyzing reviews unrelated to Gen-AI, we exclude all reviews prior to 2022, i.e., the year ChatGPT was released. While this filtering improves temporal alignment, it cannot fully guarantee that all remaining reviews pertain to post-Gen-AI functionality.
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. This single-platform scope may limit the generalizability of results to other ecosystems with different user bases or review practices. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,066 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
Our results show that LLMs can reliably extract user-discussed topics with high accuracy, i.e., 91%, when supported by five-shot prompting and data cleaning. We identify the 10 most frequently discussed topics revealing user behavior and challenges that inform actionable insights for developers and researchers. Furthermore, our findings suggest an evolution in user expectations, with a shift in user behavior, such as shifting from early adopters who were more forgiving of limitations to a mainstream user base that demands higher expectations. Future work will explore cross-platform comparisons and further investigate how expectations vary across different user profiles (e.g., casual vs. expert users) as Gen-AI tools continue and mature.
Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at omitted.
Acknowledgements.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [RGPIN-2021-03969].